A Complete Bibliography of Bell Labs Technical Journal, 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

16 March 2020 Version 1.23

Title word cross-reference

10× [Win14]. 2 [CGP^{+}12]. 3 [CHK12a, CGP^{+}12, HSKH13]. μ [KPK12].

100 [FLR^{+}10, RWS^{+}10, RBPC^{+}10]. 101 [KML13].

2.0 [FLR11, Lee10]. 25 [EGZO13].

3D [VLV11].

40 [CL10b].

50 [CL10b]. 50th [BWW14].

8 [RBPC^{+}10].

Absolute [AS12]. Access [JB12, MN15, SdLvWtB13, CC11, FLR11, GNM^{+}10, KC10, KES10, OBT^{+}11]. adaptable [VBB11]. adaptation [BEK^{+}10, FS10]. Adaptive [BER11, DDR14, HDWV12, Rad12, RJC12, RTG13, TVV^{+}13, VDB^{+}13, KET10, WC10]. addition [NGYG11]. Advanced [FBGG10, SW13]. advances [SCR^{+}10]. Advantageous [DDM^{+}13]. advertising [KC10]. Affordances [VGF13]. afforded [KET10]. age [Ger12]. agent [Yas11]. aggregation [DLK^{+}10, EZ11]. aggregators [KT11a]. agile [FBGG10]. air [Her10]. Alcatel [Ano10a, Ano10b, Ano10c, Ano10d, Ano11c, Ano11d, Ano11e, Ano11f, Ano12i, Ano12j, Ano12k, Ano12l, Ano13i, Ano13j, Ano13k, Ano13l, Ano14c, FBC14].

Algorithms [MGMP13, San15, GKW^{+}10]. allocation [CPP10, WC10, ZYG10]. American [Ger12]. amplification [BHSW10]. amplified [KCWL10]. Analysis [DMUV13, DDM^{+}13, JZS^{+}14, MJ14, PUM^{+}13, AKM^{+}10]. Analytical [PG13]. Analytics [CHV^{+}14, FBC14, HMOS12]. Anniversary [BWW14]. Annual [Kor13]. Answer [BK13]. antenna [CT11, ZYG10]. Antennas [HHtBD13]. Application [CK13, KR11, KMP^{+}13, Lee10, MJ14, Rad12, AKM^{+}10, ABJ^{+}10, GSL^{+}11, MS11, Yas11]. Applications [BSV12b, CBS12, TBC^{+}14, CT11, EINU11, FGM11, GHPP11, KCGP10, MSZB11, Mam11, MSA^{+}11, PV11, SSG^{+}11]. Approach [AMV12, BF12, CBS12, PG13, EINU11, FOS10, RKC10]. approaches [SKL10]. Architecture [KWS12, BCIG^{+}11, BDD^{+}10, HK11]. Architectures [DDM^{+}13, VPC^{+}12, Her10, THR10, Yas11]. Areas [BDL^{+}13]. ASIC [SSM^{+}13]. ASIC/FPGA [SSM^{+}13]. Ask [BK13]. Assessing [MM12]. Assessment [RJC12, URRH13, FOS10]. assignment [GKW^{+}10]. Assisted [VPC^{+}12]. Assurance [CN12]. Attacks [BSV12a, KDP^{+}12]. Audio [MMDZ13]. Authentication [BSV12b]. Author [Ano13m, Ano14d]. Author/Subject [Ano13m, Ano14d]. Auto [BMFD13]. Auto-Calibration [BMFD13]. Automatic [AS12]. Autonomous [ACFG10, AAH^{+}10, MGP10]. Availability [Rad12, ABJ^{+}10]. Aware [MHP^{+}12, CCT11].

Background [BWW14]. Backhauling [BDL^{+}13, DLK^{+}10]. Backplane [MZZ13]. Backup [KNSW12]. band [CC11]. Bandwidth [AMV12, BCIG^{+}11, WC10]. Bandwidth-Effective [AMV12]. Banking [KML13]. bankless [Mam11]. Base [DDM^{+}13, HDGL^{+}13, VDB^{+}13, CT11].

Based [CFS^{+}12, HDWV12, MMDZ13, Sha12, VVG12, BEK^{+}10, FZ10, GDFC10, HK11, KES10, MZ11, MSA^{+}11, MS11, NGYG11, TQ10]. Beamforming [HSKH13]. Behavior [HMOS12]. Bell [Ger12, Kim10, NG11]. benefits [EZ11]. between [SSG^{+}11]. Beyond [AO12, EGZO13, FLR11, Tka10]. Big [San15, STD^{+}14]. Birth [Bow12, Lal14]. Bit [LLB^{+}13, VDB^{+}13, KET10]. Bits [VGF13]. blending [BCF^{+}11]. books [FZ10]. Broadband [JB12, THU12, KES10]. Broadcast [FMKR12, MMD12, URRH13]. build [McG11]. Building [MGCSA^{+}13]. Built [McG11]. Burst [CHSW13, MZ11]. Burst-Switched [CHSW13]. bus [DLK^{+}10]. Business [Bow12, CDF12, STD^{+}14, BCIG^{+}11, DDV11, KHV11].

C [KPK12]. C/OS [KPK12]. C/OS-II [KPK12]. Cache [KPK12]. caching [DR11]. CAEMP [CCT11]. Calibration [AS12, BMFD13]. Cameras [AS12]. Capacities [Win14]. Capacity [CKK^{+}13, EFKW10, LCCV13]. Carrier [TDK^{+}19, NGYG11]. Carrier-grade [TDK^{+}19]. Case [BEF^{+}12, JB12, VVG12, GSL^{+}11]. Casensa [CCT11]. cases [BCIG^{+}11]. cathedral [Dys12]. CDN [DR11]. Cell [CKK^{+}13, DMUV13]. Cell/Wi [DMUV13]. Cell/Wi-Fi [DMUV13]. Cells [BMFD13, HHtBD13]. Cellular [BKKR13, BSV12a, Del14, HSKH13, LCCV13, Yan12, ACFG10]. Celnet [BDN^{+}10]. Centered [GG13]. Centric [BKKR13, MV12, EINU11]. Challenge [CGP^{+}12]. Challenges [CSMGAFD13, FGM11, SC12, VVG12, AB11]. Change [FMKR12]. Channel [FMKR12, KCWL10]. Channels [MMD12, RBPC^{+}10]. Characteristics [BDL^{+}13]. Characterization [KPK12]. Characterizing [BF12]. Charging [LBC12, CL10a]. choose [PV11]. Churn [PUM^{+}13]. Cliff [KFS12]. Cloud [Don12, HDGL^{+}13, BEF^{+}12, CHK12a, CFS^{+}12, CHK^{+}12b, MW12, Mul12, Rad12, Sha12]. Cloud-Based [CFS^{+}12, Sha12]. clouds [HS19]. Co [Leh13]. Co-Creation [Leh13]. Coding [JLHC^{+}12]. cognitive [KLMN10]. Coherent [LLB^{+}13, RTG13, RBPC^{+}10]. Combining [LARJ12]. come [McG11]. Common [EGZO13]. Communication [BDD^{+}10, LBW^{+}12, DKT11, LM11, Rad11, SCR^{+}10, VLV11, XWH^{+}10, Yas11]. Communications [VW14, BBI^{+}11, BCF^{+}11, MS11, Tka10, YJ11]. communities [McG11]. Community [MSA^{+}11, FGM11]. Community-based [MSA^{+}11]. compensators [XWH^{+}10]. Complex [KWS12]. Complexity [MJ14]. compliance [YJ11]. Composition [KMP^{+}13]. Compound [Kor13]. Compression [GCTS13]. Compressive [JLHC^{+}12, JZS^{+}14]. Computations [San15]. Computing [BEF^{+}12, MW12, Mul12]. concept [VLV11]. Concepts [TVV^{+}13]. Conditions [URRH13]. conferencing [dLvWEM^{+}10]. Configurable [LMV^{+}13]. Congestion [CK13]. Connected [VPC^{+}12]. connection [GFS11]. connectivity [CC11, KU11]. Connects [Zam13]. Considerations [MCP12, YJ11]. Constraint [GDFC10]. Constraint-based [GDFC10]. consumer [KHV11]. consumers [Mam11]. consumption [LH10]. contact [FZ10]. contact-books [FZ10]. Content [KWS12, Sha12, VPC^{+}12, CNAM11, JMM^{+}11, DDV11]. contention [NGYG11]. contention-based [NGYG11]. context [BHK^{+}11, CCT11]. context-aware [CCT11]. Continuity [CDF12]. Control [AU12, LBC12, BCIG^{+}11, FLR11, FBGG10, TDK^{+}19, XW11]. Controlled [TVV^{+}13]. Converged [JB12, KC10]. Convergence [PRV^{+}12]. Conversion [LT13]. cooled [Her10]. Cooperation [BKKR13, SSG^{+}11]. Copper [MN15, GNM^{+}10]. Core [SdLvWtB13, Win14, BGR11]. Correction [SdLvWtB13]. Cosmic [BWW14]. Cost [NGYG11, PG13, PDP^{+}13, KES10, JMM^{+}11]. Cover [Ano11a, Ano12d, Ano12a, Ano12b, Ano12c, Ano13a, Ano13b, Ano13c, Ano13d, Ano14a]. coverage [RKC10]. CPRI [GCTS13]. Creation [Leh13, Lee10]. critical [BCD^{+}11, LM11, Rad11]. Cross [Zam13, KC10]. Cross-Connects [Zam13]. cross-domain [KC10]. crowdsourcing [GMN^{+}11]. Cues [AS12]. curmudgeon [And11]. Current [Zam13]. Customer [CHV^{+}14, STD^{+}14, OBT^{+}11].

D [CHK12a, CGP^{+}12, HSKH13]. Data [DMUV13, FBC14, MMDZ13, NDZ13, San15, STD^{+}14, ESMUVZ11, KT11b, MSZB11, MZ11]. Datacenter [CDF12]. DBPSK [CL10b]. decade [Tka10]. Decoding [FMKR12]. Deconstructing [CCT11]. Defect [AO12]. Delay [FMKR12]. Delivered [DDR14]. Delivering [BDM^{+}10, CN12]. Delivery [AMV12, CR13, MHP^{+}12, JMM^{+}11, KT11b, ZYG10]. Demand [MCHY14]. Deployed [THU12]. Deployment [BMFD13, DMUV13]. Description [Kor13]. Design [AO12, CKK^{+}13, GG13, LLB^{+}13, Leh13, SSM^{+}13, VGF13, BDD^{+}10, LH10]. Detecting [BSV12a]. Detection [Don12, LLB^{+}13, AAH^{+}10, MCS11, RBPC^{+}10]. Determination [DDM^{+}13]. developer [McG11]. device [SKL10]. devices [GFS11, JMM^{+}11, MCS11]. diagnostic [BHSW10]. didn’t [McG11]. different [And11]. Differential [Yan12]. Differentiated [DKT11]. Difficulty [KNSW12]. digital [Dys12, Liu10, MGP10, SKPW17]. Discovery [BWW14, EINU11]. discussion [DDV11]. dispersion [RBPC^{+}10, XWH^{+}10]. Distance [RTG13]. Distance-Adaptive [RTG13]. Distributed [APWA12, GKW^{+}10, GDFC10]. distribution [BHK^{+}11, CNAM11, HDD11]. diversity [CT11]. Division [FB13, PDP^{+}13, LBTC10]. DL [GKW^{+}10]. DMME [APWA12]. DO [NGYG11]. domain [KC10, TDK^{+}19]. doped [KCWL10]. DQPSK [CL10b, RWS^{+}10]. Drivers [GKD^{+}13]. DRM [MV12]. DTV [CGP^{+}12]. Dual [DLK^{+}10]. During [MPR13]. DVB [FMKR12]. DWDM [CL10b]. Dynamic [CHV^{+}14, CAH10, KDP^{+}12, LBC12, San15, CPP10, ESMUVZ11, FLR11, MGP10, WC10].

E-band [CC11]. E-mobility [BHK^{+}11]. easy [GMN^{+}11]. easy-to-use [GMN^{+}11]. Eco [THR10, MNPW10, MMN^{+}10]. eco-friendly [MNPW10]. Eco-sustainable [THR10, MMN^{+}10]. economically [GHPP11]. Economies [KWS12]. edge [HS19]. Effect [KFS12, PMRG11]. Effective [AMV12, BMFD13, NGYG11]. efficiency [FS10, GN11, Her10, KCGP10, KLMN10]. Efficient [GKD^{+}13, SSM^{+}13, DLK^{+}10, EBD^{+}10, FBB^{+}10, GMN^{+}11, GNM^{+}10, MMN^{+}10, PCSS11]. eHealth [SC12]. Elastic [CHK^{+}12b, LMV^{+}13]. electric [BHK^{+}11]. Electrical [EGZO13]. Embedded [MCS11]. Emerging [BSV12b, KML13, GMN^{+}11]. EMI [YJ11]. Empirical [Kor13]. empowering [CCT11]. enable [FBB^{+}10]. enabled [KC10]. Enablement [CBS12, KMP^{+}13, AKM^{+}10, GSL^{+}11, KR11, MS11, Yas11]. enabler [GDFC10]. Enablers [HDD11, Lee10]. Enabling [PCSS11, GMN^{+}11]. encoding [TQ10]. Encryption [KS12]. End [ABJ^{+}10, CBS12, XW11]. End-to-End [CBS12, ABJ^{+}10]. endpoint [MCS11]. Energy [BEK^{+}10, EBD^{+}10, KPK12, KCGP10, LH10, SSM^{+}13, BHK^{+}11, BAP10, FBB^{+}10, FS10, GN11, Her10, MMN^{+}10]. Energy-efficient [EBD^{+}10, FBB^{+}10, MMN^{+}10]. energy-saving [BAP10]. Engine [TBC^{+}14]. Engineering [URRH13]. Enhanced [Her10, KLMN10, MGCSA^{+}13]. Enhancements [MPS^{+}13]. Enterprise [BF12, FBC14, BCF^{+}11, FHLV10]. Entities [KB13]. Entity [APWA12]. environment [CNAM11]. environmental [FOS10, YJ11]. Environments [CSMGAFD13]. EPC [CL10a]. EPC/LTE [CL10a]. Equipment [MZZ13, Her10, LH10, YJ11]. era [FLR^{+}10, SKPW17]. erbium [KCWL10]. Error [SdLvWtB13]. Estimating [PG13]. EV [NGYG11]. EV-DO [NGYG11]. Evaluation [GN11, MGMP13]. event [FGM11]. Evolution [BDL^{+}13, CSMGAFD13, KWS12, LMV^{+}13, GS10]. Evolved [URRH13]. examples [CPP10]. Excess [HHtBD13]. execution [CNAM11]. Expansion [JB12]. Expectations [KET10]. Experience [BMT13, CHV^{+}14, CGP^{+}12, GMRH13, PL13, STD^{+}14, BDM^{+}10, NGYG11, OBT^{+}11, PMRG11]. Exploiting [PG13]. Exploring [ABKN14]. Explosion [DMUV13, ESMUVZ11]. extended [GS10]. extensions [DLK^{+}10, MS11]. extraction [SKL10].

Facial [TQ10]. factory [Ger12]. Fade [Yan12]. Fast [FMKR12]. Fault [Rad12]. Fault-Tolerance [Rad12]. Feasibility [CSMGAFD13]. Feature [BWW14, SKL10]. FEC [FMKR12]. Feeds [MMDZ13]. Femtocell [MGCSA^{+}13]. femtocells [ACFG10, CAH10, SKL10]. Fi [DMUV13]. Fiber [Win14, EFKW10, KCWL10, LBTC10]. Fibers [FB13, RBPC^{+}10]. field [RWS^{+}10]. Financial [KML13]. Fingerprint [MGCSA^{+}13]. Fingerprinting [MMWH14, SKL10]. firewall [BGR11]. Fixed [JB12, KHV11]. Flexible [PG13, RTG13, Zam13]. Flexible-Grid [PG13]. Flow [MJ14, SSM^{+}13]. Forecasting [MCHY14]. formats [CL10b]. Forum [EGZO13]. Forward [SdLvWtB13]. Fostering [SSG^{+}11]. FPGA [SSM^{+}13]. fractal [TQ10]. fragmented [Mam11]. Framework [KMP^{+}13, LBC12]. Free [MZZ13]. Free-Space [MZZ13]. Frequency [MMWH14, CC11]. Freudian [Liu10]. friendly [MNPW10]. Front [Ano11a, Ano12d, Ano12a, Ano12b, Ano12c, Ano13a, Ano13b, Ano13c, Ano13d, Ano14a]. FTTH [KES10]. FTTH-based [KES10]. Fuel [BES13]. Functional [KPK12]. Functional-Level [KPK12]. Fusing [AS12]. Future [CGDS^{+}12, HS19, MN15, VW14, Yas11, Zam13, EBD^{+}10, KR11, Liu10, THR10]. fuzzy [RKC10].

Gain [Yan12, KET10]. Gateways [THU12]. Gb [CL10b, RWS^{+}10, RBPC^{+}10, EGZO13, FLR^{+}10]. Gb/s [CL10b, RWS^{+}10, RBPC^{+}10, EGZO13, FLR^{+}10]. General [PV11, MS11]. Generation [BDL^{+}13, BES13, SVHM12, CSS^{+}10, CC11]. geofencing [GHPP11]. Gesture [MGMP13]. GHz [CL10b, CC11]. Global [LMV^{+}13]. government [BBI^{+}11]. GPON [WC10]. grade [TDK^{+}19]. Graded [FB13]. Graded-Index [FB13]. great [Ger12]. green [MNPW10]. Greening [MPS^{+}13]. Grid [KT11b, MM12, PG13, RTG13, DKT11, YJ11]. Grids [MCHY14, BHK^{+}11, BDD^{+}10]. Group [BSV12b, FB13]. Growth [BES13, Kor13, LCCV13, KHV11, SKPW17]. Guaranteed [VDB^{+}13, PV11].

Hacking [Bow12]. Handoff [Yan12]. handover [AAH^{+}10]. handsets [NdLvW11]. haul [BHSW10, RBPC^{+}10]. HDTV [MPR13]. Headend [PRV^{+}12]. healthcare [MSZB11]. help [BER11].

Heterogeneous [LCCV13, SW13, OBT^{+}11]. High [CSMGAFD13, CFS^{+}12, GKD^{+}13, LLB^{+}13, LT13, MSZB11, XWH^{+}10, CKK^{+}13]. high-speed [XWH^{+}10]. Hijacking [BSV12a]. HSPA [LMMT11]. HTTP [DDR14, HDWV12, RJC12, VDB^{+}13]. Human [PL13]. Hyperbolic [Kor13].

ICIC [GKW^{+}10]. idea [Ger12]. identification [SKL10]. identity [FHLV10]. idle [CAH10]. II [KPK12]. image [TQ10]. Impact [ABKN14, KHV11, NDZ13, FOS10]. impairments [LBTC10]. Improve [STD^{+}14]. Improvement [HSKH13, NGYG11]. Improving [OBT^{+}11, Rad12]. IMS [Lee10]. In-Building [MGCSA^{+}13]. Inclusion [KML13]. increasing [FOS10]. Index [Ano13m, Ano14d, FB13]. India [KML13]. Indoor [BMFD13]. Industrial [VVG12]. Industry [Bow12]. Information [Ano11b, Ano12h, Ano12e, Ano12f, Ano12g, Ano13e, Ano13f, Ano13g, Ano13h, Ano14b, MJ14, EFKW10]. Inherent [KNSW12]. Initiation [BGR11]. Innovation [LV13, Ger12, NG11]. Integrated [ELN^{+}14]. Integrative [AO12]. Intelligent [GFS11]. Interaction [Leh13, VGF13, CBL^{+}11]. Interactions [MGMP13]. Interactive [AMV12, Sha12]. Interconnection [CDF12]. Interconnects [EGZO13]. Interface [EGZO13, GCTS13, VGF13]. Interference [BKKR13, LARJ12, DGG^{+}10]. intermediated [CNAM11]. internet [DTV^{+}10, EBD^{+}10, EZ11, Kor13, KWS12]. Internetworking [EGZO13]. interworking [GFS11]. Intra [EGZO13]. Intra-System [EGZO13]. Intrinsically [SVHM12]. Introduction [Mar15, VV12, WM11, Her10]. Intrusion [Don12]. Investigating [BK13]. IP [BGR11, Del14, HK11, KU11]. IP-based [HK11]. Issue [Ano11b, Ano12h, Ano12e, Ano12f, Ano12g, Ano13e, Ano13f, Ano13g, Ano13h, Ano14b]. ITU [BES13]. izing [KML13].

Join [MMDZ13]. Joint [MJ14, ZYG10]. journal [Lal14].

Key [KES10]. Knowledge [BK13].

laboratory [BBI^{+}11]. Labs [Kim10, Ger12, NG11]. LambdaXtreme(R) [RWS^{+}10]. Large [San15]. Latency [CFS^{+}12, JZS^{+}14, Sha12]. Latent [BF12]. Layer [CHV^{+}14, PG13, MKVD10, TDK^{+}19]. learning [RKC10]. letter [Kim10]. Level [GKD^{+}13, KPK12]. Leverage [THU12]. Leveraging [JB12]. LGS [BBI^{+}11]. lifecycle [FOS10]. lightRadio^{T M} [CKK^{+}13]. Likelihood [KDP^{+}12]. Limits [KS12, EFKW10]. line [MGP10]. linear [HDD11]. Live [RWS^{+}10]. load [MZ11]. Localization [MMWH14]. Location [CR13, MS11, XW11]. Locking [KPK12]. Long [CSMGAFD13, BHSW10, RBPC^{+}10]. loss [PMRG11]. Low [FMKR12, JZS^{+}14, PDP^{+}13, LH10, RBPC^{+}10, JMM^{+}11]. Low-cost [JMM^{+}11]. LSH [ABKN14]. LTE [APWA12, AAH^{+}10, BCD^{+}11, BDN^{+}10, CL10a, CK13, CR13, DDR14, DGG^{+}10, DDM^{+}13, GKW^{+}10, GCTS13, LARJ12, MHP^{+}12, RKC10, SW13, URRH13]. LTE-Advanced [SW13]. LTE/LTE [GCTS13]. Lucent [Ano10a, Ano10b, Ano10c, Ano10d, Ano11c, Ano11d, Ano11e, Ano11f, Ano12i, Ano12j, Ano12k, Ano12l, Ano13i, Ano13j, Ano13k, Ano13l, Ano14c, FBC14].

Machine [BSV12a]. Machine-to-Machine [BSV12a]. Macro [DMUV13, Lee10]. macro-enablers [Lee10]. maintain [McG11]. Making [HHtBD13]. malware [MCS11]. Management [APWA12, BKKR13, CGDS^{+}12, KDP^{+}12, MMDZ13, CPP10, DGG^{+}10, FHLV10, FBGG10, GDFC10, MNPW10, NdLvW11]. manager [GFS11]. Managing [KU11]. mango [JMM^{+}11]. Manycore [BEF^{+}12]. Margin [Yan12]. market [GMN^{+}11, GHPP11, KT11a, Mam11]. markets [AB11, MSZB11]. mass [GHPP11]. mass-market [GHPP11]. Massive [HHtBD13, Mar15]. Measuring [Del14, NDZ13]. mechanism [ACCV10]. mechanisms [FHLV10]. Media [KB13, Liu10]. Mediated [LBW^{+}12]. meeting [FHLV10]. memoirs [NG11]. Method [FMKR12, LT13, PV11, ZML10]. metrics [GN11]. Metro [CKK^{+}13, RTG13, Win14]. Metro/Core [Win14]. microgrid [KT11a]. Microwave [BWW14, BDL^{+}13]. Migration [EGZO13, JB12]. MIMO [CT11, HHtBD13, Mar15, ZYG10]. Mission [LM11, BCD^{+}11, Rad11]. Mitigating [CFS^{+}12]. Mitigation [LBTC10]. Mobile [AS12, BMT13, CNAM11, DMUV13, ESMUVZ11, EINU11, FMKR12, KML13, MHP^{+}12, NDZ13, THU12, VDB^{+}13, VW14, BEK^{+}10, FZ10, FBB^{+}10, FGM11, GMN^{+}11, GFS11, JMM^{+}11, KHV11, LMMT11, NdLvW11]. Mobile-izing [KML13]. Mobility [APWA12, BHK^{+}11]. Mode [FB13, CAH10, XWH^{+}10]. Model [CHV^{+}14, DDR14, GG13, KT11a, KHV11, VBB11]. Modeling [HMOS12, MNPW10]. Models [AO12, VVG12, BCIG^{+}11, GSL^{+}11, DDV11]. Modern [Lal14]. Modulation [TVV^{+}13]. Monetizing [ESMUVZ11]. MoneyBee [GMN^{+}11]. Monitoring [MSZB11]. MPLS [ACCV10]. MPLS-TP [ACCV10]. Multi [CHV^{+}14, GKD^{+}13, PG13, PRV^{+}12, dLvWEM^{+}10, GFS11, KU11, TDK^{+}19]. multi-connectivity [KU11]. multi-domain [TDK^{+}19]. Multi-Layer [CHV^{+}14, PG13, TDK^{+}19]. Multi-Level [GKD^{+}13]. Multi-stream [dLvWEM^{+}10]. multi-technology [GFS11]. Multi-Vendor [PRV^{+}12]. multicarrier [BAP10]. Multicast [URRH13]. Multidimensional [BF12]. Multidisciplinary [GMRH13]. Multimedia [BGR11, MMDZ13, URRH13, BDM^{+}10]. Multimode [FB13]. Multiple [AS12]. multiplexed [LBTC10]. Multiplexing [FB13, PDP^{+}13, Win14]. multisource [AKM^{+}10].

Needs [KB13, FHLV10]. neighbor [AAH^{+}10]. Network [BKKR13, BMT13, CN12, DDR14, MV12, MCP12, MHP^{+}12, PG13, PUM^{+}13, SSM^{+}13, THU12, URRH13, VPC^{+}12, BER11, BCIG^{+}11, BDD^{+}10, CL10a, CC11, EZ11, FZ10, FBGG10, GKW^{+}10, KET10, KHV11, KLMN10, KES10, LMMT11, MNPW10, PV11, RWS^{+}10, THR10, TDK^{+}19, dLvWEM^{+}10]. Network-Centric [BKKR13, MV12]. Networking [FBC14, LV13, VDD12, BCF^{+}11, FLR^{+}10, HC10]. Networks [BKKR13, BSV12a, BF12, CHSW13, CGDS^{+}12, DMUV13, HSKH13, LMV^{+}13, LARJ12, LCCV13, MJ14, MGCSA^{+}13, MPS^{+}13, RTG13, SVHM12, San15, SdLvWtB13, SW13, Yan12, ACCV10, ACFG10, BDM^{+}10, BHSW10, BEK^{+}10, BDN^{+}10, CPP10, CSS^{+}10, DKT11, DLK^{+}10, EFKW10, FBGG10, FS10, GN11, KCGP10, KHV11, LM11, MGP10, MMN^{+}10, MKVD10, Rad11, RKC10, SSG^{+}11, THR10, TDK^{+}19, ZML10]. Neuron [GSL^{+}11]. Next [BDL^{+}13, BES13, CC11, SVHM12, CSS^{+}10, Tka10]. Next-Generation [BDL^{+}13, BES13, SVHM12, CC11]. NIX [BEF^{+}12]. NoBot [MCS11]. node [KC10]. Nodes [SSM^{+}13]. Non [GG13, CNAM11, HDD11]. non-intermediated [CNAM11]. non-linear [HDD11]. Non-Place [GG13]. nonlinear [LBTC10]. notes [And11].

Novel [CBS12, Her10].

OADMs [CSS^{+}10]. OFDM

[LBTC10, ZYG10]. Offload [THU12]. Offloading [NDZ13]. Omnidirectional [AMV12]. Ongoing [KDP^{+}12]. Online [CL10a, MV12]. Open [BCIG^{+}11]. Operating [MNPW10, RWS^{+}10]. Operations [BMT13, Rad11]. operator [FHLV10]. Opportunities [MW12, SC12, AB11, KR11]. opportunity [ESMUVZ11]. Optical [CHSW13, EGZO13, GKD^{+}13, HC10, LMV^{+}13, LV13, MZZ13, MPS^{+}13, PG13, RTG13, SdLvWtB13, XWH^{+}10, DLK^{+}10, FBGG10, FLR^{+}10, HK11, KET10, LH10, MKVD10, SCR^{+}10, Tka10, ZML10]. optically [EFKW10]. optically-routed [EFKW10]. Optics [Win14]. Optimal [MPR13, PG13]. Optimization [MMD12, TBC^{+}14, VDB^{+}13, AAH^{+}10, BDN^{+}10, CPP10, MGP10, RKC10]. Optimized [CT11, SSM^{+}13, KES10]. Optimizing [VDD12]. Optimum [DR11]. Options [DMUV13]. Order [KS12]. Order-Preserving [KS12]. organized [GKW^{+}10, KLMN10]. organizing [DGG^{+}10]. origins [Dys12]. OS-II [KPK12]. Outliers [CFS^{+}12]. overlay [SSG^{+}11]. Overload [AU12]. Overview [BS10, DP12, HL13, HJ14, Koz10, LW11, SB10, UV13, DDV11].

packaging [FOS10]. Packet

[BDL^{+}13, CSS^{+}10, FS10, PMRG11]. paired [RBPC^{+}10]. Paradigm [BSV12b]. Parallel [TBC^{+}14]. Parameters [ABKN14, KES10]. Past [MN15, VW14]. Patents [Ano12i, Ano12j, Ano12k, Ano12l, Ano13i, Ano13j, Ano13k, Ano13l, Ano14c, Ano10a, Ano10b, Ano10c, Ano10d, Ano11c, Ano11d, Ano11e, Ano11f]. path [KR11]. patterns [CBL^{+}11]. Payload [LT13]. PDM [RBPC^{+}10]. PDM-QPSK [RBPC^{+}10]. peer [dLvWEM^{+}10]. peer-to-peer [dLvWEM^{+}10]. Peering [KWS12]. Performance [CK13, HSKH13, LT13, STD^{+}14, URRH13, MSZB11]. Period [MPR13]. periodic [KT11b]. personal [DTV^{+}10]. Personalization [ABKN14, KC10]. personalization-enabled [KC10]. Personalized [AKM^{+}10, VBB11]. Perspective [CHSW13]. Petabit [MZZ13]. Petabit/s [MZZ13]. photonic [FBGG10]. Photonics [ELN^{+}14]. Physics [Lal14]. pipes [ESMUVZ11]. Place [GG13]. Place-Centered [GG13]. plan [ZML10]. plane [FBGG10]. Planning [MMN^{+}10, GDFC10, KES10]. platform [CCT11]. Playout [FMKR12]. polarization [CT11, LBTC10, XWH^{+}10]. policies [ESMUVZ11]. Policy [LBC12]. Positioning [BMFD13, MGCSA^{+}13]. potential [DLK^{+}10]. Power [FBB^{+}10, GNM^{+}10, KCWL10, KT11a, LH10, NdLvW11, ZYG10]. Power-efficient [GNM^{+}10]. Practical [MMD12]. Predictable [Mul12]. Prediction [AO12, MZ11, PUM^{+}13]. Prediction-based [MZ11]. Prefatory [Kim10]. Preliminary [VGF13]. PresenceScape [LBW^{+}12]. Present [MN15, VW14]. Preserving [ABKN14, KS12]. President [Kim10]. Preventing [BSV12a]. Primary [KNSW12]. Primary-Backup [KNSW12]. principles [BDD^{+}10]. Privacy [ABKN14, KS12, XW11]. Privacy-Preserving [ABKN14]. Probabilistic [MMWH14]. procedures [CAH10]. Process [VVG12]. Processing [DDM^{+}13]. productivity [SKPW17]. profiling [ACCV10, AKM^{+}10]. Programming [VGF13]. Project [GMRH13, GSL^{+}11]. Projections [Kor13]. Protection [VPC^{+}12]. Protocol [MPS^{+}13, BGR11, KT11b]. Prototype [MZZ13]. provide [CC11]. Provided [KS12]. Providers [MW12, PV11]. Providing [CGP^{+}12]. PSTN [JB12]. Public [BCD^{+}11, HMOS12, MCP12, MSZB11, Rad11].

QoE [DDR14]. QoS [ACCV10, DKT11, ESMUVZ11, EZ11, PV11]. QPSK [RBPC^{+}10]. Quality [CGP^{+}12, KFS12, RJC12, BDM^{+}10, BCIG^{+}11, KHV11, PMRG11]. Quantifying [EZ11]. Quantitative [DDM^{+}13]. Qubit [SML18]. Questions [BK13]. Quo [SML18].

Radio [BDL^{+}13, HDGL^{+}13, MMWH14, URRH13, CC11, FBB^{+}10]. Radio-Frequency [MMWH14]. Railway [CSMGAFD13]. railways [LM11]. Raman [BHSW10]. Range [MZZ13]. Rapid [AU12]. Rate [FS10, Kor13, LLB^{+}13, VDB^{+}13, KET10]. Real [CBS12, VLV11, KT11b]. Real-time [VLV11]. Realistic [URRH13, ZML10]. realization [VLV11]. Realizing [CPP10, FLR^{+}10]. recognition [TQ10]. Recognizing [FHLV10]. recommendation [DDV11]. Recommender [VDD12]. reconfigurable [KCWL10]. reconfiguration [KLMN10]. recovery [MKVD10]. Redirecting [NDZ13]. Reducing [FOS10, San15]. reinforcement [RKC10]. Reinvent [GMRH13]. Rejection [LARJ12]. relation [AAH^{+}10]. Relationships [BF12]. Relative [AS12]. Reliability [AO12, FOS10, Her10]. Reliable [CN12, KT11b, PCSS11]. remittance [Mam11]. Rendering [CHK12a]. Replication [KNSW12]. Reproduction [KFS12]. Research [GMRH13, BBI^{+}11]. resource [CPP10, GKW^{+}10, KLMN10]. Result [VGF13]. Results [TVV^{+}13]. return [SKPW17]. Revenue [GSL^{+}11]. Reviews [Lal14]. RF [SKL10]. Rich [LBW^{+}12, JMM^{+}11]. right [McG11]. Ring [RTG13, CSS^{+}10]. Risk [KDP^{+}12]. roaming [CL10a]. robot [Liu10]. ROCS [AU12]. Role [VGF13, Yas11]. routed [EFKW10]. Routing [Zam13]. Run [MMWH14].

s [CL10b, MZZ13, RWS^{+}10, RBPC^{+}10, EGZO13, FLR^{+}10, RBPC^{+}10]. Safety [MCP12, BCD^{+}11, MSZB11, Rad11, YJ11]. Saving [KPK12, BAP10]. savings [BEK^{+}10]. Scalable [CHK^{+}12b, JLHC^{+}12, San15, ZYG10, JMM^{+}11]. Scaling [Rad12, Tka10, Win14]. scavenging [KCGP10]. Scenarios [BDL^{+}13, MV12]. scene [Lal14]. SCF [KMP^{+}13]. Schemes [BKKR13]. Science [PL13]. scientific [Lal14]. scoring [KC10]. SDN [TDK^{+}19]. SDTV [MPR13]. seamless [GFS11]. Secure [CBS12, CN12, Mam11, SVHM12, KT11b]. Securing [BES13]. Security [CDF12, MM12, MCP12, SC12, FHLV10, HK11]. selection [ZYG10]. Self [BDN^{+}10, DGG^{+}10, WC10, CPP10, GKW^{+}10, GDFC10, KLMN10, RKC10]. Self-adaptive [WC10]. self-management [CPP10, GDFC10]. Self-optimization [BDN^{+}10, CPP10, RKC10]. self-organized [GKW^{+}10, KLMN10]. Self-organizing [DGG^{+}10]. Semi [Kor13]. Semi-Empirical [Kor13]. Sensing [JLHC^{+}12, JZS^{+}14]. Sensor [HMOS12]. Sequence [MJ14]. Service [KMP^{+}13, URRH13, ABJ^{+}10, BCIG^{+}11, CBL^{+}11, FZ10, GMN^{+}11, MS11, XW11]. Services [CFS^{+}12, CHK^{+}12b, CGP^{+}12, GG13, Leh13, MMD12, BCD^{+}11, DKT11, EZ11, KU11, KR11, Yas11]. session [AKM^{+}10, BGR11]. Set [MPR13, PDP^{+}13]. SGTP [KT11b]. SH [FMKR12]. sharing [ACFG10, XW11]. shedding [MZ11]. Signal [GCTS13]. Similarity [MMDZ13]. Simulating [Del14]. Simulation [AU12]. Situations [MMD12]. Size [MPR13].

SlideWorld [GMRH13]. Slowdown [FMKR12]. Small [DMUV13, HHtBD13, San15]. Smart [HHtBD13, MM12, NdLvW11, YJ11, BDD^{+}10, DKT11, KT11b, MCHY14]. Social [BF12, FBC14, KB13, MJ14, PUM^{+}13, BCF^{+}11, FZ10]. Software [AO12, LMV^{+}13, Yas11]. Solution [DMUV13, PRV^{+}12, FZ10, GHPP11]. Solutions [CN12, CR13]. SON [GS10]. Space [HMOS12, MZZ13]. Spatial [Win14, CT11]. Special [BWW14]. Spectrally [GKD^{+}13]. Spectrum [BES13, ACFG10]. Speed [CSMGAFD13, GKD^{+}13, XWH^{+}10]. Split [DDM^{+}13]. Stadiums [CKK^{+}13]. stage [KT11a]. Standardization [EGZO13, BAP10]. Start [FMKR12]. Station [DDM^{+}13, VDB^{+}13, CT11]. Stations [HDGL^{+}13]. statistics [BEK^{+}10]. Storage [VDD12]. Strategic [GG13]. strategies [DR11]. Stream [MMDZ13, dLvWEM^{+}10]. Streaming [DDR14, FMKR12, HDWV12, RJC12, Sha12, TBC^{+}14, VDB^{+}13, BER11]. streams [MSZB11, MZ11]. Study [JB12, VVG12, GSL^{+}11]. Subject [Ano13m, Ano14d]. Subjective [RJC12]. submarine [SCR^{+}10]. Subscriber [PUM^{+}13, EINU11, FZ10, MGP10]. subscriber-centric [EINU11]. Subsystem [BGR11]. Success [KDP^{+}12]. Support [BDL^{+}13, LMMT11, ABJ^{+}10, KC10, Rad11]. Suppressing [KFS12]. Surveillance [JZS^{+}14]. sustainable [MMN^{+}10, THR10]. SVC [HDWV12]. SVC-Based [HDWV12]. Switched [CHSW13]. Switching [Del14]. synchronized [FGM11]. System [BEF^{+}12, Del14, EGZO13, LLB^{+}13, MMDZ13, URRH13, THR10]. Systems [Don12, ELN^{+}14, FMKR12, GKD^{+}13, LLB^{+}13, VDD12, FBB^{+}10, HC10, HK11, KCWL10, Lee10, Rad11, SCR^{+}10, XWH^{+}10, ZYG10]. takes [Lal14]. Tangible [VGF13]. Tb [RBPC^{+}10]. Tb/s [RBPC^{+}10]. TDD [HHtBD13]. technical [BCIG^{+}11, VLV11]. Technologies [BES13, CR13, RJC12, SW13, FLR11, KES10, Lee10, DDV11]. Technology [LV13, PG13, GFS11]. Telco [TBC^{+}14, DR11]. Telecom [MW12, FHLV10, MNPW10]. Telecommunication [CFS^{+}12, CHK^{+}12b, MZZ13, Her10, VBB11]. Telecommunications [AB11, MMN^{+}10]. Term [CSMGAFD13]. Terrestrial [LLB^{+}13]. testing [YJ11]. their [BAP10]. Theory [ABJ^{+}10, MSZB11]. thermal [Her10]. Thermally [ELN^{+}14]. Things [CBL^{+}11]. thoughtful [And11]. Time [CBS12, KT11b, VLV11]. Timely [KNSW12]. Tolerance [Rad12]. Tool [AU12, LLB^{+}13, BHSW10]. TP [ACCV10]. Traffic [Del14, Kor13, NDZ13, THU12, BEK^{+}10, KET10, KHV11]. transaction [Mam11]. transactions [KT11a]. Transceivers [RTG13, BAP10, FBB^{+}10]. Transcoding [MHP^{+}12]. Transforming [BCF^{+}11]. transients [KCWL10]. Transition [MPR13]. Transmission [GKD^{+}13, LLB^{+}13, MMD12, CL10b, EFKW10, HK11, KCWL10, KET10, LBTC10, RWS^{+}10, RBPC^{+}10]. Transmitter [TVV^{+}13]. Transmitter-Controlled [TVV^{+}13]. transparent [CL10b, MKVD10, ZML10]. Transport [KT11b, ACCV10, EBD^{+}10, FLR^{+}10, THR10]. Trends [Kor13, FLR11]. trial [RWS^{+}10]. Trust [CGDS^{+}12]. Trusted [THU12]. Turing [Dys12]. TV [LMMT11]. TWDM [PDP^{+}13]. two [KT11a, CPP10]. two-stage [KT11a].

U.S. [BBI^{+}11]. ubiquitous [GMN^{+}11]. UL [GKW^{+}10]. ultra [BHSW10]. ultra-long-haul [BHSW10]. unconscious [Liu10]. underlying [SSG^{+}11]. Understanding [KB13]. Unicast [MMD12]. unified [BCF^{+}11, XW11]. Unifying [CGDS^{+}12]. universe [Dys12]. unstructured [ACFG10]. Uplink [BKKR13, PCSS11]. Use [HHtBD13, GMN^{+}11, BCIG^{+}11]. User [BMT13, GMRH13, KB13, MGMP13, PL13, AKM^{+}10, NGYG11, XW11]. Using [DDR14, JLHC^{+}12, JZS^{+}14, Kor13, MHP^{+}12, MGCSA^{+}13, PUM^{+}13, STD^{+}14, THU12, CC11, RBPC^{+}10, ZYG10]. Utilizing [KC10, BDN^{+}10].

Vadis [SML18]. Validation [VVG12]. Values [Leh13]. Variable [FMKR12]. Vendor [PRV^{+}12]. vertical [AB11, MSZB11]. via [CPP10, KT11a]. viable [GHPP11]. Video [AMV12, CGP^{+}12, DDR14, FMKR12, HMOS12, JLHC^{+}12, JZS^{+}14, KFS12, MGMP13, MHP^{+}12, PRV^{+}12, RJC12, VDD12, VDB^{+}13, DDV11, EZ11, HDD11, PCSS11, PMRG11, RWS^{+}10, VLV11, ZYG10, dLvWEM^{+}10]. Videoconferencing [GMRH13]. Virtual [Don12, LBW^{+}12, VLV11]. Vision [ELN^{+}14, CBL^{+}11, FLR^{+}10]. Visual [AS12]. Visualization [KB13]. Visualizing [BF12]. voice [BCD^{+}11]. voices [NG11]. Volume [Ano13m, Ano14d].

Wavelength [MKVD10, PDP^{+}13, Zam13]. Wavelength-Set [PDP^{+}13]. way [McG11]. Web [AKM^{+}10, CBL^{+}11, FLR11, Lee10, VBB11]. Wi-Fi [DMUV13]. wideband [CC11]. Will [SKPW17]. WiMAX [CT11]. Wireless [BMFD13, BDL^{+}13, BES13, Yan12, CC11, GN11, KLMN10, OBT^{+}11, PCSS11, ZYG10]. Within [FBC14, KB13]. WLAN [THU12]. Workers [BK13]. World [LBW^{+}12, VPC^{+}12]. worlds [VLV11].

Xplorer [BDN^{+}10].

References

Aziz:2010:ANR

[AAH^{+}10] Danish Aziz, Anton Ambrosy, Lester T. W. Ho, Lutz Ewe, Markus Gruber, and Hajo Bakker. Autonomous neighbor relation detection and handover optimization in LTE. Bell Labs Technical Journal, 15(3):63–83, December 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Acharya:2011:TVM

[AB11] Swarup Acharya and Kenneth C. Budka. Telecommunications in vertical markets: Challenges and opportunities. Bell Labs Technical Journal, 16(3):1–4, December 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Asthana:2010:EES

[ABJ^{+}10] Abhaya Asthana, Robert S.

Blake, Richard D. Jordan, Reza Shafie-Khorasani, and Daryl J. Steen. End-to-end service availability support: Theory and application. Bell Labs Technical Journal, 15 (1):149–173, June 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Aghasaryan:2014:EIL

[ABKN14] Armen Aghasaryan, Makram Bouzid, Dimitre Kostadinov, and Animesh Nandi. Exploring the impact of LSH parameters in privacy-preserving personalization. Bell Labs Technical Journal, 18(4): 33–44, March 2014. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Addeo:2010:QMP

[ACCV10] Christian Addeo, Giorgio Cazzaniga, Roberta Crescentini, and Fabio Valente. On QoS mechanism profiling in MPLS-TP transport networks. Bell Labs Technical Journal, 14(4):227–241, Winter 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Andrews:2010:ASS

[ACFG10] Matthew Andrews, Véronique Capdevielle, Afef Feki, and Piyush Gupta. Autonomous spectrum sharing for unstructured cellular networks with femtocells. Bell Labs Technical Journal, 15(3):85– 97, December 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Aghasaryan:2010:PAE

[AKM^{+}10] Armen Aghasaryan, Murali Kodialam, Sarit Mukherjee, Yann Toms, Christophe Senot, Stéphane Betgé-Brezetz, T. V. Lakshman, and Limin Wang. Personalized application enablement by Web session analysis and multisource user profiling. Bell Labs Technical Journal, 15(1):67–76, June 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Alface:2012:IOV

[AMV12] Patrice Rondao Alface, Jean-François Macq, and Nico Verzijp. Interactive omnidirectional video delivery: a bandwidth-effective approach. Bell Labs Technical Journal, 16(4):135– 147, March 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anderson:2011:MDN

[And11] Philip W. Anderson. More and different: notes from a thoughtful curmudgeon. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2011. ISBN 981-4350-12-5 (hardcover), 981-4350-13-3 (paperback). 2011 pp. LCCN Q171 .A527 2011.

Anonymous:2010:RALa

[Ano10a] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 14(4):329– 340, Winter 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2010:RALb

[Ano10b] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 15(1): 237–258, June 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2010:RALc

[Ano10c] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 15(2):245– 266, September 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2010:RALd

[Ano10d] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 15(3):207– 232, December 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2011:FC

[Ano11a] Anonymous. Front cover. Bell Labs Technical Journal, 16(3):i, December 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2011:II

[Ano11b] Anonymous. Issue information. Bell Labs Technical Journal, 16(3):ii, December 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2011:RALa

[Ano11c] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 15(4):189– 212, March 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2011:RALb

[Ano11d] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 16(1): 235–258, June 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2011:RALc

[Ano11e] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 16(2):277– 298, September 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2011:RALd

[Ano11f] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 16(3):181– 211, December 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2012:FCa

[Ano12a] Anonymous. Front cover.

Bell Labs Technical Journal, 16(4):??, March 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2012:FCb

[Ano12b] Anonymous. Front cover. Bell Labs Technical Journal, 17(1):??, June 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2012:FCc

[Ano12c] Anonymous. Front cover. Bell Labs Technical Journal, 17(2):??, September 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2012:FC

[Ano12d] Anonymous. Front cover. Bell Labs Technical Journal, 17(3):??, December 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2012:IIa

[Ano12e]  Anonymous.  Issue infor-

mation. Bell Labs Technical Journal, 16(4):??, March 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2012:IIb

[Ano12f] Anonymous. Issue information. Bell Labs Technical Journal, 17(1):??, June 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2012:IIc

[Ano12g] Anonymous. Issue information. Bell Labs Technical Journal, 17(2):??, September 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2012:II

[Ano12h] Anonymous. Issue information. Bell Labs Technical Journal, 17(3):??, December 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2012:RALa

[Ano12i] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 16(4):243– 267, March 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2012:RALb

[Ano12j] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 17(1): 247–274, June 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2012:RALc

[Ano12k] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 17(2):143– 171, September 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2012:RALd

[Ano12l] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 17(3):213– 239, December 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2013:FCa

[Ano13a] Anonymous. Front cover. Bell Labs Technical Journal, 17(4):??, March 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2013:FCb

[Ano13b] Anonymous. Front cover. Bell Labs Technical Journal, 18(1), June 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2013:FCc

[Ano13c] Anonymous. Front cover. Bell Labs Technical Journal, 18(2), September 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2013:FCd

[Ano13d] Anonymous. Front cover. Bell Labs Technical Journal, 18(3):c1, December 2013. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2013:IIa

[Ano13e]  Anonymous.  Issue infor-

mation. Bell Labs Technical Journal, 17(4):??, March 2013. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2013:IIb

[Ano13f] Anonymous. Issue information. Bell Labs Technical Journal, 18(1), June 2013. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2013:IIc

[Ano13g]  Anonymous.  Issue infor-

mation. Bell Labs Technical Journal, 18(2), September 2013. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2013:IId

[Ano13h] Anonymous. Issue information. Bell Labs Technical Journal, 18(3):??, December 2013. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2013:RALa

[Ano13i] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 17(4):175– 207, March 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2013:RALb

[Ano13j] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 18(1): 213–244, June 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2013:RALc

[Ano13k] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 18(2):255– 287, September 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2013:RALd

[Ano13l] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 18(3):293– 321, December 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2013:VAS

[Ano13m] Anonymous. Volume 17 author/subject index. Bell Labs Technical Journal, 17(4):209– 217, March 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2014:FC

[Ano14a] Anonymous. Front cover. Bell Labs Technical Journal, 18(4):??, March 2014. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2014:II

[Ano14b]  Anonymous.  Issue infor-

mation. Bell Labs Technical Journal, 18(4):??, March 2014. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2014:RAL

[Ano14c] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 18(4):199– 230, March 2014. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2014:VAS

[Ano14d] Anonymous. Volume 18 author/subject index. Bell Labs Technical Journal, 18(4):231– 243, March 2014. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Asthana:2012:ISD

[AO12] Abhaya Asthana and Kazu Okumoto. Integrative software design for reliability: Beyond models and defect prediction. Bell Labs Technical Journal, 17(3):37– 59, December 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

An:2012:DDL

[APWA12] Xueli An, Fabio Pianese, Indra Widjaja, and Utku Günay Acer. DMME: a distributed LTE mobility management entity. Bell Labs Technical Journal, 17(2):97– 120, September 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Aerts:2012:ACM

[AS12] Maarten Aerts and Erwin Six. Automatic calibration for mobile cameras by fusing multiple relative and absolute visual cues. Bell Labs Technical Journal, 16(4):187– 202, March 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Akyamac:2012:RTR

[AU12] Ahmet A. Akyamac and Huseyin Uzunalioglu. ROCS: a tool for rapid overload control simulation. Bell Labs Technical Journal, 17 (1):67–87, June 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bitzer:2010:NES

[BAP10] Thomas Bitzer, Jacques Achard, and Andreas Pascht. New energy-saving multicarrier transceivers and their standardization. Bell Labs Technical Journal, 15(2):53– 58, September 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Beacken:2011:LGC

[BBI^{+}11] Marc Beacken, Linda Braun, Dominick J. Imbesi, Lloyd G. Greenwald, Michael J. Geller, Adrian Hartman, Michael J. Chen, Mike Ter Louw, Alfonso Piccirilli, and David Bishop. LGS’ government communications laboratory and research for the U.S. government. Bell Labs Technical Journal, 16(3):5–28, December 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Budka:2011:PSM

[BCD^{+}11] Kenneth C. Budka, Thomas Chu, Tewfik L. Doumi, Wim Brouwer, Phil Lamoureux, and Maria E. Palamara. Public safety mission critical voice services over LTE. Bell Labs Technical Journal, 16 (3):133–149, December 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Burns:2011:TEC

[BCF^{+}11] Michael J. Burns, R. Bruce Craig Jr., Brian D. Friedman, Peter D. Schott, and Christophe Senot. Transforming enterprise communications through the blending of social networking and unified communications. Bell Labs Technical Journal, 16 (1):19–34, June 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Brunetti:2011:ONQ

[BCIG^{+}11] James A. Brunetti, Kanti Chakrabarti, Alina M. Ionescu-Graff, Ramesh Nagarajan, and Dong Sun. Open network quality of service and bandwidth control: Use cases, technical architecture, and business models. Bell Labs Technical Journal, 16(2):133– 152, September 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Budka:2010:CNA

[BDD^{+}10] Kenneth C. Budka, Jayant G. Deshpande, Tewfik L. Doumi, Mark Madden, and Tim Mew. Communication network architecture and design principles for smart grids. Bell Labs Technical Journal, 15(2):205–227, September 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Boiocchi:2013:NGM

[BDL^{+}13] Gianluca Boiocchi, Paolo Di Prisco, Ahmed Lahrech, Pierre Lopez, Maurizio Moretto, and Paolo Volpato. Nextgeneration microwave packet radio: Characteristics and evolution areas to support new scenarios in wireless backhauling. Bell Labs Technical Journal, 18(2):143– 157, September 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Batteram:2010:DQE

[BDM^{+}10] Harold Batteram, Gerard Damm, Amit Mukhopadhyay, Laurent Philippart, Rhodo Odysseos, and Carlos Urrutia-Valdés. Delivering quality of experience in multimedia networks. Bell Labs Technical Journal, 15 (1):175–193, June 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Buvaneswari:2010:SOL

[BDN^{+}10]  Arumugam  Buvaneswari,

Lawrence Drabeck, Nachi Nithi, Mark Haner, Paul Polakos, and Chitra Sawkar. Self-optimization of LTE networks utilizing Celnet Xplorer. Bell Labs Technical Journal, 15(3):99–117, December 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ballesteros:2012:NCM

[BEF^{+}12] Francisco J. Ballesteros, Noah Evans, Charles Forsyth, Gorka Guardiola, Jim McKie, Ron Minnich, and Enrique Soriano-Salvador. NIX: a case for a manycore system for Cloud computing. Bell Labs Technical Journal, 17(2):41– 54, September 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Blume:2010:ESM

[BEK^{+}10] Oliver Blume, Harald Eckhardt, Siegfried Klein, Edgar Kuehn, and Wieslawa M. Wajda. Energy savings in mobile networks based on adaptation to traffic statistics. Bell Labs Technical Journal, 15(2):77– 94, September 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Benno:2011:ASN

[BER11] Steven Benno, Jairo O. Esteban, and Ivica Rimac. Adaptive streaming: The network HAS to help. Bell Labs Technical Journal, 16(2):101– 114, September 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bonin:2013:SST

[BES13] Jean-Pierre Bonin, Cengiz Evci, and Amy L. Sanders. Securing spectrum through the ITU to fuel the growth of next-generation wireless technologies. Bell Labs Technical Journal, 18(2):99– 115, September 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Burns:2012:MAC

[BF12] Michael J. Burns and Brian D. Friedman. A multidimensional approach to characterizing and visualizing latent relationships in enterprise social networks. Bell Labs Technical Journal, 17 (1):201–217, June 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bessis:2011:SIP

[BGR11] Thierry Bessis, Vijay K. Gurbani, and Ashwin Rana. Session Initiation Protocol firewall for the IP Multimedia Subsystem core. Bell Labs Technical Journal, 15(4):169– 187, March 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ballentin:2011:MCE

[BHK^{+}11] Ralph Ballentin, Detlef Hartmann, Dieter Kopp, Thomas Loewel, Matthias Sund, and Wolfgang Templ. E-mobility in the context of electric energy distribution grids. Bell Labs Technical Journal, 16 (3):47–60, December 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bengtsson:2010:DTR

[BHSW10] Thomas Bengtsson, Tin Kam Ho, Todd Salamon, and Christopher A. White. A diagnostic tool for Raman amplification in ultra-long-haul networks. Bell Labs Technical Journal, 14(4):61–72, Winter 2010. CODEN BLTJFD.

ISSN 1089-7089 (print), 1538-7305 (electronic).

Burns:2013:QAQ

[BK13] Michael J. Burns and Xerxes P. Kotval. Questions about questions: Investigating how knowledge workers ask and answer questions. Bell Labs Technical Journal, 17(4): 43–61, March 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Balachandran:2013:NCC

[BKKR13] Krishna Balachandran, Joseph H. Kang, Kemal M. Karakayali, and Kiran M. Rege. Networkcentric cooperation schemes for uplink interference management in cellular networks. Bell Labs Technical Journal, 18(2):23–36, September 2013. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Barba:2013:WIP

[BMFD13] Félix Barba Barba, Aaron Garrido Martin, and Alfonso Fernández-Durán. Wireless indoor positioning: Effective deployment of cells and auto-calibration. Bell Labs Technical Journal, 18(2):213– 235, September 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Barshefsky:2013:MUE

[BMT13] Alvin Barshefsky, Mohamad M.

Mehio, and Hemi Trickey. Mobile user experience for network operations. Bell Labs Technical Journal, 17 (4):5–20, March 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bowles:2012:BHB

[Bow12] Matt Bowles. The business of hacking and birth of an industry. Bell Labs Technical Journal, 17(3):5–16, December 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Benowitz:2010:O

[BS10] Marc Benowitz and Sam Samuel. Overview. Bell Labs Technical Journal, 15(2):1– 5, September 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Broustis:2012:DPM

[BSV12a] Ioannis Broustis, Ganapathy S. Sundaram, and Harish Viswanathan. Detecting and preventing machineto-machine hijacking attacks in cellular networks. Bell Labs Technical Journal, 17 (1):125–140, June 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Broustis:2012:GAN

[BSV12b] Ioannis Broustis, Ganapathy S. Sundaram, and Harish Viswanathan. Group authentication: a new paradigm for emerging applications. Bell Labs Technical Journal, 17 (3):157–173, December 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bahr:2014:ASF

[BWW14] C. Bahr, M. Weldon, and R. W. Wilson. 50th anniversary special feature: The discovery of cosmic microwave background. Bell Labs Technical Journal, 19(1):1–7, ???? 2014. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Claussen:2010:DIM

[CAH10] Holger Claussen, Imran Ashraf, and Lester T. W. Ho. Dynamic idle mode procedures for femtocells. Bell Labs Technical Journal, 15(2):95– 116, September 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Christophe:2011:WTV

[CBL^{+}11] Benoit Christophe, Mathieu Boussard, Monique Lu, Alain Pastor, and Vincent Toubiana. The Web of Things vision: Things as a service and interaction patterns. Bell Labs Technical Journal, 16 (1):55–61, June 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Cakulev:2012:SER

[CBS12] Violeta Cakulev, Ioannis Broustis, and Ganapathy S. Sundaram. Secure enablement of real time applications: a novel end-to-end approach. Bell Labs Technical Journal, 17(3):175– 191, December 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Colombo:2011:NGA

[CC11] Claudio Colombo and Massimo Cirigliano. Nextgeneration access network: a wireless network using Eband radio frequency (71– 86 GHz) to provide wideband connectivity. Bell Labs Technical Journal, 16 (1):187–205, June 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Criel:2011:DCC

[CCT11] Johan Criel, Laurence Claeys, and Lieven Trappeniers. Deconstructing Casensa: The CAEMP context-aware empowering platform. Bell Labs Technical Journal, 16 (1):35–53, June 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ceballos:2012:BCS

[CDF12] Juan Ceballos, Richard Di-Pasquale, and Robert Feldman. Business continuity and security in datacenter interconnection. Bell Labs Technical Journal, 17(3):147– 155, December 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chang:2012:MHL

[CFS^{+}12] Fangzhe Chang, Peter S. Fales, Moritz Steiner, Ramesh Viswanathan, Thomas J. Williams, and Thomas L. Wood. Mitigating high latency outliers for Cloudbased telecommunication services. Bell Labs Technical Journal, 17(2):121– 142, September 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ciavaglia:2012:UMF

[CGDS^{+}12] Laurent Ciavaglia, Samir Ghamri-Doudane, Mikhail Smirnov, Panagiotis Demestichas, Vera-Alexandra Stavroulaki, Aimilia Bantouna, and Berna Sayrac. Unifying management of future networks with trust. Bell Labs Technical Journal, 17 (3):193–212, December 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Cubero:2012:PVS

[CGP^{+}12] José María Cubero, Jesús Gutiérrez, Pablo Pérez, Enrique Estalayo, Julián Cabrera, Fernando Jaureguizar, and Narciso García. Providing 3D video services: The challenge from 2D to 3DTV quality of experience. Bell Labs Technical Journal, 16(4):115–134, March 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Carroll:2012:RC

[CHK12a] Martin D. Carroll, Ilija Hadžić, and William A. Katsak. 3D rendering in the Cloud. Bell Labs Technical Journal, 17(2):55–66, September 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chang:2012:SET

[CHK^{+}12b] Yuh-Jye Chang, Adiseshu Hari, Pramod Koppol, Antony Martin, and Thanos Stathopoulos. Scalable and elastic telecommunication services in the Cloud. Bell Labs Technical Journal, 17(2):81– 96, September 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Cazzaniga:2013:NPB

[CHSW13] Giorgio Cazzaniga, Christian Hermsmeyer, Iraj Saniee, and Indra Widjaja. A new perspective on burst-switched optical networks. Bell Labs Technical Journal, 18(3):111– 131, December 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chen:2014:MLD

[CHV^{+}14] Sining Chen, Tin Kam Ho, Avinash Vyas, Jin Cao, and Jeffrey Spiess. A multi-layer dynamic model for customer experience analytics. Bell Labs Technical Journal, 18 (4):19–32, March 2014. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Calin:2013:LAC

[CK13] Doru Calin and Bong Ho Kim. LTE application and congestion performance. Bell Labs Technical Journal, 18 (1):5–25, June 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Calin:2013:HCL

[CKK^{+}13] Doru Calin, Aliye Özge Kaya, Bong Ho Kim, Kai Yang, and Simon Yiu. On the High Capacity lightRadio^{T M} metro cell design for stadiums. Bell Labs Technical Journal, 18 (2):77–97, September 2013. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Cai:2010:OCR

[CL10a] Yigang Cai and Xiang Yang Li. Online charging in the roaming EPC/LTE network. Bell Labs Technical Journal, 15(1):115–132, June 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chandrasekhar:2010:GDD

[CL10b] Sethumadhavan Chandrasekhar and Xiang Liu. 40 Gb/s DBPSK and DQPSK formats for transparent 50 GHz DWDM transmission. Bell Labs Technical Journal, 14 (4):11–25, Winter 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chandrashekhar:2012:DNA

[CN12] Uma Chandrashekhar and Scott Nelson. Delivering network assurance through secure and reliable solutions. Bell Labs Technical Journal, 17(3):1–4, December 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Christophe:2011:MEE

[CNAM11] Benoit Christophe, Maribel Narganes, Ville Antila, and Linas Maknavicius. Mobile execution environment for non-intermediated content distribution. Bell Labs Technical Journal, 15(4):117– 134, March 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Carofiglio:2010:RSM

[CPP10] Giovanna Carofiglio, Pierre Peloso, and Hélia Pouyllau. Realizing self-management via self-optimization in dynamic networks: Two examples of dynamic resource allocation. Bell Labs Technical Journal, 15(3):177– 192, December 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Cherian:2013:LLT

[CR13] Suma S. Cherian and Ashok N. Rudrapatna. LTE location technologies and delivery solutions. Bell Labs Technical Journal, 18(2):175– 194, September 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Calle-Sanchez:2013:LTE

[CSMGAFD13] Jaime Calle-Sánchez, Mariano Molina-García, José I. Alonso, and Alfonso Fernández-Durán. Long term evolution in high speed railway environments: Feasibility and challenges. Bell Labs Technical Journal, 18(2):237– 253, September 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chiaroni:2010:PON

[CSS^{+}10] Dominique Chiaroni, Géma Buforn Santamaria, Christian Simonneau, Sophie Etienne, Jean-Christophe Antona, Sébastien Bigo, and Jesse Simsarian. Packet OADMs for the next generation of ring networks. Bell Labs Technical Journal, 14 (4):265–283, Winter 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Charaabi:2011:OWM

[CT11] Zied Charaabi and Marc Testard. Optimized WiMAX MIMO antenna for base station applications with polarization and spatial diversity. Bell Labs Technical Journal, 16(1):217–234, June 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dotsch:2013:QAS

[DDM^{+}13] Uwe Dötsch, Mark Doll, Hans-Peter Mayer, Frank Schaich, Jonathan Segel, and Philippe Sehier. Quantitative analysis of split base station processing and determination of advantageous architectures for LTE. Bell Labs Technical Journal, 18 (1):105–128, June 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

DeVriendt:2014:QMV

[DDR14] Johan De Vriendt, Danny De Vleeschauwer, and Dave C. Robinson. QoE model for video delivered over an LTE network using HTTP adaptive streaming. Bell Labs Technical Journal, 18(4): 45–62, March 2014. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Vriendt:2011:VCR

[DDV11] Johan De Vriendt, Natalie Degrande, and Marc Verhoeyen. Video content recommendation: an overview and discussion on technologies and business models. Bell Labs Technical Journal, 16 (2):235–250, September 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

DelSignore:2014:MSC

[Del14] Kenneth W. Del Signore. Measuring and simulating cellular switching system IP traffic. Bell Labs Technical Journal, 18(4):159– 180, March 2014. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dominique:2010:SOI

[DGG^{+}10] Francis Dominique, Christian G. Gerlach, Nandu Gopalakrishnan, Anil Rao, James P. Seymour, Robert Soni, Aleksandr Stolyar, Harish Viswanathan, Carl Weaver, and Andreas Weber. Self-organizing interference management for LTE. Bell Labs Technical Journal, 15(3):19–42, December 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Deshpande:2011:DSQ

[DKT11] Jayant G. Deshpande, Eunyoung Kim, and Marina Thottan. Differentiated services QoS in smart grid communication networks. Bell Labs Technical Journal, 16(3):61– 81, December 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dorgeuille:2010:DOB

[DLK^{+}10] François Dorgeuille, Nicolas Le Sauze, Benaïssa El Khattar, Laurent Ciavaglia, Emmanuel Dotaro, and Thierry Zami. Dual optical bus for efficient aggregation and backhauling networks and potential extensions. Bell Labs Technical Journal, 14(4):243– 264, Winter 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

vanWijngaarden:2010:MSV

[dLvWEM^{+}10] Adriaan J. de Lind van Wijngaarden, Bilgehan Erman, Elissa P. Matthews, Ron Sharp, and Edward Sutter. Multi-stream video conferencing over a peer-to-peer network. Bell Labs Technical Journal, 15(2):229– 243, September 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dhawan:2013:MSC

[DMUV13] Perminder Dhawan, Amit Mukhopadhyay, and Carlos Urrutia-Valdés. Macro and small cell/wi-fi networks: An analysis of deployment options as the solution for the mobile data explosion. Bell Labs Technical Journal, 18 (1):59–79, June 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Donadio:2012:VID

[Don12] Pasquale Donadio. Virtual intrusion detection systems in the cloud. Bell Labs Technical Journal, 17(3):113– 128, December 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dinu:2012:O

[DP12] Mihaela Dinu and Michael Peeters. Overview. Bell Labs Technical Journal, 17 (1):1–4, June 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

DeVleeschauwer:2011:OCS

[DR11] Danny De Vleeschauwer and Dave C. Robinson. Optimum caching strategies for a telco CDN. Bell Labs Technical Journal, 16(2):115– 132, September 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Daenen:2010:PI

[DTV^{+}10] Koen Daenen, Bart Theeten, David Vanderfeesten, Bart Vrancken, Eric Waegeman, Jan Moons, Mihai Dragan, Monica Bojin, Mircea Strugaru, Knarig Arabshian, and Peter Bosch. The personal internet. Bell Labs Technical Journal, 15(1):3–21, June 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dyson:2012:TCO

[Dys12]  George Dyson.  Turing’s

cathedral: the origins of the digital universe. Pantheon Books, New York, NY, USA, 2012. ISBN 0-375-42277-3 (hardcover). xxii + 401 pp. LCCN QA76.17 .D97 2012.

Eilenberger:2010:EET

[EBD^{+}10] Gert J. Eilenberger, Stephan Bunse, Lars Dembeck, Ulrich Gebhard, Frank Ilchmann, Wolfram Lautenschlaeger, and Jens Milbrandt. Energyefficient transport for the future Internet. Bell Labs Technical Journal, 15(2):147– 167, September 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Essiambre:2010:CLI

[EFKW10] René-Jean Essiambre, Gerard J. Foschini, Gerhard Kramer, and Peter J. Winzer. Capacity limits of information transmission in opticallyrouted fiber networks. Bell Labs Technical Journal, 14 (4):149–162, Winter 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Eckart:2013:MOI

[EGZO13] Thomas Eckart, Hans-Joachim Götz, Thomas Ziselsberger, and Klaus-Holger Otto. The migration of the optical internetworking forum common electrical interface standardization to optical intrasystem interconnects beyond 25 Gb/s. Bell Labs Technical Journal, 18(3):267– 283, December 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Erman:2011:MAD

[EINU11] Bilgehan Erman, Ali Inan, Ramesh Nagarajan, and Huseyin Uzunalioglu. Mobile applications discovery: a subscriber-centric approach. Bell Labs Technical Journal, 15(4):135–148, March 2011.

CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Enright:2014:VTI

[ELN^{+}14] R. Enright, Shenghui Lei, K. Nolan, I. Mathews, A. Shen, G. Levaufre, R. Frizzell, Guang-Hua Duan, and D. Hernon. A vision for thermally integrated photonics systems. Bell Labs Technical Journal, 19(1):31–45, ???? 2014. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

El-Sayed:2011:MDE

[ESMUVZ11] Mohamed El-Sayed, Amit Mukhopadhyay, Carlos Urrutia-Valdés, and Z. John Zhao. Mobile data explosion: Monetizing the opportunity through dynamic policies and QoS pipes. Bell Labs Technical Journal, 16(2):79–99, September 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Erman:2011:QQB

[EZ11] Bilgehan Erman and Z. John Zhao. Quantifying QoS benefits in the aggregation network for Internet video services. Bell Labs Technical Journal, 16(2):63– 77, September 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Franz:2013:MGD

[FB13] Bernd Franz and Henning Bülow. Mode group division multiplexing in graded-index multimode fibers. Bell Labs Technical Journal, 18(3):153– 172, December 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ferling:2010:PET

[FBB^{+}10] Dieter Ferling, Thomas Bitzer, Thomas Bohn, Dirk Wiegner, and Andreas Pascht. Power efficient transceivers to enable energy-efficient mobile radio systems. Bell Labs Technical Journal, 15(2):59– 76, September 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Friedman:2014:ESN

[FBC14] Brian D. Friedman, Michael J. Burns, and Jin Cao. Enterprise social networking data analytics within Alcatel– Lucent. Bell Labs Technical Journal, 18(4):89–109, March 2014. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Fogliata:2010:ANM

[FBGG10] Paolo Fogliata, Ermanno Belloli, Pietro V. Grandi, and Matteo Gumier. Advanced network management of the optical control plane in agile photonic networks. Bell Labs Technical Journal, 15 (1):195–208, June 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Florkey:2011:CMS

[FGM11] Cynthia K. Florkey, Ruth S.

Gayde, and Todd C. Morgan. Challenges for mobile synchronized community event applications. Bell Labs Technical Journal, 15(4):111– 116, March 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Faynberg:2010:NSM

[FHLV10] Igor Faynberg, Mark A. Hartman, Hui-Lan Lu, and Douglas W. Varney. On new security mechanisms for identity management: Recognizing and meeting telecom operator and enterprise needs. Bell Labs Technical Journal, 15(1):95–113, June 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Frigerio:2010:ROT

[FLR^{+}10] Silvano Frigerio, Alberto Lometti, Juergen Rahn, Stephen Trowbridge, and Eve L. Varma. Realizing the optical transport networking vision in the 100 Gb/s era. Bell Labs Technical Journal, 14(4):163–192, Winter 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Faynberg:2011:DAC

[FLR11] Igor Faynberg, Hui-Lan Lu, and Herbert Ristock. On dynamic access control in Web 2.0 and beyond: Trends and technologies. Bell Labs Technical Journal, 16(2):199– 218, September 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Faucheux:2012:VFD

[FMKR12] Frédéric Faucheux, Marie-Line Alberi Morel, Sylvaine Kerboeuf, and Laurent Roullet. Variable FEC decoding delay and playout slowdown method for low start delay and fast channel change for video streaming in DVB– SH mobile broadcast systems. Bell Labs Technical Journal, 16(4):63–83, March 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Franey:2010:REI

[FOS10] John P. Franey, Thomas A. Okrasinski, and William J. Schaeffer Sr. Reducing environmental impact and increasing reliability through packaging: a lifecycle assessment approach. Bell Labs Technical Journal, 15(2):193– 204, September 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Francini:2010:RAE

[FS10] Andrea Francini and Dimitrios Stiliadis. Rate adaptation for energy efficiency in packet networks. Bell Labs Technical Journal, 15(2):131– 146, September 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Feng:2010:SNS

[FZ10] Xiang Feng and Qing A. Zhu.

A social network service solution based on mobile subscriber contact-books. Bell Labs Technical Journal, 15 (1):53–66, June 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Guo:2013:LLS

[GCTS13] Bin Guo, Wei Cao, An Tao, and Dragan Samardzija. LTE/LTE — a signal compression on the CPRI interface. Bell Labs Technical Journal, 18(2):117– 133, September 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ghamri-Doudane:2010:CBD

[GDFC10] Samir Ghamri-Doudane, Eric Fabre, and Laurent Ciavaglia. Constraint-based distributed planning as an enabler for self-management. Bell Labs Technical Journal, 15(3):193– 198, December 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gernter:2012:IFB

[Ger12] Jon Gernter. The idea factory: the Bell Labs and the great age of American innovation. Penguin, New York, NY, USA, 2012. ISBN 1-59420-328-8. 422 pp. LCCN TK5102.3.U6 G47 2012.

Grinshpun:2011:ICM

[GFS11] Edward Grinshpun, David W. Faucher, and Sameer Sharma. Intelligent connection manager for seamless interworking of multi-technology mobile devices. Bell Labs Technical Journal, 15(4): 5–21, March 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gaste:2013:PNP

[GG13] Yann Gasté and Annie Gentes. “place” and “nonplace”: a model for the strategic design of placecentered services. Bell Labs Technical Journal, 17(4): 21–36, March 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Greenwald:2011:EVS

[GHPP11] Amie Greenwald, Georg Hampel, Chitra Phadke, and Viswanath Poosala. An economically viable solution to geofencing for mass-market applications. Bell Labs Technical Journal, 16(2):21– 38, September 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Godin:2013:HSM

[GKD^{+}13] Jean Godin, Agnieszka Kończykowska, Jean-Yves Dupuy, Muriel Riet, Virginie Nodjiadjim, Filipe Jorge, Gabriel Charlet, Oriol Bertran-Pardo, Jérémie Renaudier, Haik Mardoyan, Alan Gnauck, and Peter J. Winzer. High speed multilevel drivers for spectrally efficient optical transmission systems. Bell Labs Technical Journal, 18(3):67–94, December 2013. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gerlach:2010:IDU

[GKW^{+}10] Christian G. Gerlach, Ingo Karla, Andreas Weber, Lutz Ewe, Hajo Bakker, Edgar Kuehn, and Anil Rao. ICIC in DL and UL with network distributed and self-organized resource assignment algorithms in LTE. Bell Labs Technical Journal, 15(3):43– 62, December 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Govindaraj:2011:MTE

[GMN^{+}11] Dinesh Govindaraj, Naidu K. V. M., Animesh Nandi, Girija Narlikar, and Viswanath Poosala. MoneyBee: Towards enabling a ubiquitous, efficient, and easy-to-use mobile crowdsourcing service in the emerging market. Bell Labs Technical Journal, 15 (4):79–92, March 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gonguet:2013:SMR

[GMRH13] Arnaud Gonguet, Olivier Martinot, Florentin Rodio, and Vincent Hiribarren. SlideWorld: a multidisciplinary research project to reinvent the videoconferencing user experience. Bell Labs Technical Journal, 17(4):133– 144, March 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gandhi:2011:EEE

[GN11] Asif D. Gandhi and Mark E. Newbury. Evaluation of the energy efficiency metrics for wireless networks. Bell Labs Technical Journal, 16 (1):207–215, June 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Guenach:2010:PEC

[GNM^{+}10] Mamoun Guenach, Carl Nuzman, Jochen Maes, Michael Peeters, Yan Li, Danny Van Bruyssel, and Frank Defoort. Power-efficient copper access. Bell Labs Technical Journal, 15(2):117–129, September 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Graybeal:2010:ESE

[GS10] John M. Graybeal and Kamakshi Sridhar. The evolution of SON to extended SON. Bell Labs Technical Journal, 15(3):5–18, December 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gollakota:2011:RMA

[GSL^{+}11] Ram Gollakota, Abdol Saleh, Anne Lee, Veena Mendiratta, Daisy Su, and Guy J. Zenner. Revenue models for application enablement: The Project Neuron case study. Bell Labs Technical Journal, 16(2):263– 269, September 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hamaide:2010:OSN

[HC10] Jean-Pierre Hamaide and Andrew Chraplyvy. Optical systems and networking. Bell Labs Technical Journal, 14 (4):1–2, Winter 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Huysegems:2011:ENL

[HDD11] Raf Huysegems, Bart De Vleeschauwer, and Koen De Schepper. Enablers for nonlinear video distribution. Bell Labs Technical Journal, 16 (1):77–90, June 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Haberland:2013:RBS

[HDGL^{+}13] Bernd Haberland, Fariborz Derakhshan, Heidrun Grob-Lipski, Ralf Klotsche, Werner Rehm, Peter Schefczik, and Michael Soellner. Radio base stations in the cloud. Bell Labs Technical Journal, 18 (1):129–152, June 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Huysegems:2012:SBH

[HDWV12] Raf Huysegems, Bart De Vleeschauwer, Tingyao Wu, and Werner Van Leekwijck. SVC-based HTTP adaptive streaming. Bell Labs Technical Journal, 16(4): 25–41, March 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hernon:2010:EEE

[Her10] Domhnaill Hernon. Enhanced energy efficiency and reliability of telecommunication equipment with the introduction of novel air cooled thermal architectures. Bell Labs Technical Journal, 15(2):31– 51, September 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hoydis:2013:MSU

[HHtBD13] Jakob Hoydis, Kianoush Hosseini, Stephan ten Brink, and Mérouane Debbah. Making smart use of excess antennas: Massive MIMO, small cells, and TDD. Bell Labs Technical Journal, 18(2):5– 21, September 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ho:2014:O

[HJ14] Tin Kam Ho and Philippe Jacquet. Overview. Bell Labs Technical Journal, 18 (4):1–2, March 2014. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hofmann:2011:TSA

[HK11] Stefan Hofmann and Rudolf Kasseckert. Towards a security architecture for IP-based optical transmission systems. Bell Labs Technical Journal, 16(1):133–153, June 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hilt:2013:O

[HL13] Volker Hilt and Catherine Lefèvre. Overview. Bell Labs Technical Journal, 18 (1):1–3, June 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ho:2012:PSB

[HMOS12] Tin Kam Ho, Kim Matthews, Lawrence O’Gorman, and Harald Steck. Public space behavior modeling with video and sensor analytics. Bell Labs Technical Journal, 16 (4):203–217, March 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hilt:2019:FEC

[HS19] Volker Hilt and Kevin Sparks. Future edge clouds. Bell Labs Technical Journal, 24: 1–17, December 2019. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Halbauer:2013:BPI

[HSKH13] Hardy Halbauer, Stephan Saur, Johannes Koppenborg, and Cornelis Hoek. 3D beamforming: Performance improvement for cellular networks. Bell Labs Technical Journal, 18(2):37– 56, September 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Joseph:2012:LBA

[JB12] Jean-Philippe Joseph and William J. Barnett Jr. Leveraging broadband access expansion for PSTN migration: a case study for converged fixed access. Bell Labs Technical Journal, 17 (1):141–162, June 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Jiang:2012:SVC

[JLHC^{+}12] Hong Jiang, Chengbo Li, Raziel Haimi-Cohen, Paul A. Wilford, and Yin Zhang. Scalable video coding using compressive sensing. Bell Labs Technical Journal, 16 (4):149–169, March 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Jaiswal:2011:MLC

[JMM^{+}11] Sharad Jaiswal, Anirban Majumder, Naidu K. V. M., Girija Narlikar, Viswanath Poosala, Nisheeth Shrivastava, and Ankur Jain. mango: Low-cost, scalable delivery of rich content on mobile devices. Bell Labs Technical Journal, 15(4):63–78, March 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Jiang:2014:SVA

[JZS^{+}14] Hong Jiang, Songqing Zhao, Zuowei Shen, Wei Deng, Paul A. Wilford, and Raziel Haimi-Cohen. Surveillance video analysis using compressive sensing with low latency. Bell Labs Technical Journal, 18(4):63–74, March 2014. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kotval:2013:VEW

[KB13] Xerxes P. Kotval and Michael J. Burns. Visualization of entities within social media: Toward understanding users’ needs. Bell Labs Technical Journal, 17(4):77–101, March 2013. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kisel:2010:UPE

[KC10] Andrey Kisel and Mark M. Clougherty. Utilizing a personalization-enabled access node in support of converged cross-domain scoring and advertising. Bell Labs Technical Journal, 15 (1):77–93, June 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kim:2010:ESE

[KCGP10] Kyoung Joon Kim, Francesco Cottone, Suresh Goyal, and Jeff Punch. Energy scavenging for energy efficiency in networks and applications. Bell Labs Technical Journal, 15(2):7–29, September 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kilper:2010:CPT

[KCWL10] Daniel C. Kilper, Sethumadhavan Chandrasekhar, Christopher A. White, and Bruno Lavigne. Channel power transients in erbium doped fiber amplified reconfigurable transmission systems. Bell Labs Technical Journal, 14(4):73–84, Winter 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kanoun:2012:TDR

[KDP^{+}12] Wael Kanoun, Samuel Dubus, Serge Papillon, Nora Cuppens-Boulahia, and Frédéric Cuppens. Towards dynamic risk management: Success likelihood of ongoing attacks. Bell Labs Technical Journal, 17 (3):61–78, December 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kulkarni:2010:FBB

[KES10] Samrat Kulkarni and Mohamed El-Sayed. FTTHbased broadband access technologies: Key parameters for cost optimized network planning. Bell Labs Technical Journal, 14(4):297–309, Winter 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Korotky:2010:EON

[KET10] Steven K. Korotky, René-Jean Essiambre, and Robert W.

Tkach. Expectations of optical network traffic gain afforded by bit rate adaptive transmission. Bell Labs Technical Journal, 14(4):285– 295, Winter 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kokalj-Filipovic:2012:SCE

[KFS12] Silvija Kokalj-Filipović and Emina Soljanin. Suppressing the cliff effect in video reproduction quality. Bell Labs Technical Journal, 16(4):171– 185, March 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Krogfoss:2011:ICT

[KHV11] Bill Krogfoss, Gary Hanson, and R. J. Vale. Impact of consumer traffic growth on mobile and fixed networks: Business model and network quality impact. Bell Labs Technical Journal, 16 (1):105–120, June 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kim:2010:PLB

[Kim10] Jeong Kim. Prefatory letter from Bell Labs President. Bell Labs Technical Journal, 15(2):i, September 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kuehn:2010:EWN

[KLMN10] Edgar Kuehn, Thomas Loewel, Geneviève Mange, and Klaus Nolte. Enhanced wireless network efficiency by cognitive, self-organized resource reconfiguration. Bell Labs Technical Journal, 15(3):199– 206, December 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kumar:2013:BMI

[KML13] Neha Kumar, Akhil Mathur, and Siddhartha Lal. Banking 101: Mobile-izing financial inclusion in an emerging India. Bell Labs Technical Journal, 17(4):37–41, March 2013. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kocan:2013:SFA

[KMP^{+}13] Kristin F. Kocan, Kevin W.

McKiou, Ninoo R. Pandya, Jorge Hernandez-Herrero, and F. Theodore Freuler. The SCF: A framework for application enablement service composition. Bell Labs Technical Journal, 18(1): 171–193, June 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Koppol:2012:IDT

[KNSW12] Pramod Koppol, Kedar S.

Namjoshi, Thanos Stathopoulos, and Gordon T. Wilfong. The inherent difficulty of timely primarybackup replication. Bell Labs Technical Journal, 17(2):15– 24, September 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Korotky:2013:SED

[Kor13] Steven K. Korotky. Semiempirical description and projections of Internet traffic trends using a hyperbolic compound annual growth rate. Bell Labs Technical Journal, 18(3):5–21, December 2013. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kozik:2010:O

[Koz10] Jack Kozik. Overview. Bell Labs Technical Journal, 15 (1):1–2, June 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kang:2012:FLE

[KPK12] Kyungtae Kang, Kyung-Joon Park, and Hongseok Kim. Functional-level energy characterization of μC/OS-II and cache locking for energy saving. Bell Labs Technical Journal, 17(1):219–227, June 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kocan:2011:AEP

[KR11] Kristin F. Kocan and Dave C.

Robinson. Application enablement: a path to new opportunities for future services. Bell Labs Technical Journal, 16(2):1–4, September 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kolesnikov:2012:LPP

[KS12] Vladimir Kolesnikov and Abdullatif Shikfa. On the limits of privacy provided by orderpreserving encryption. Bell Labs Technical Journal, 17 (3):135–146, December 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kim:2011:TSM

[KT11a] Hongseok Kim and Marina Thottan. A two-stage market model for microgrid power transactions via aggregators. Bell Labs Technical Journal, 16(3):101–107, December 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kim:2011:SSG

[KT11b] Young-Jin Kim and Marina Thottan. SGTP: Smart Grid Transport Protocol for secure reliable delivery of periodic real time data. Bell Labs Technical Journal, 16(3):83– 99, December 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kahn:2011:MMC

[KU11] Colin Kahn and Alistair Urie. Managing multi-connectivity for IP services. Bell Labs Technical Journal, 15(4): 45–62, March 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Krogfoss:2012:IAE

[KWS12] Bill Krogfoss, Marcus Weldon, and Lev Sofman. Internet architecture evolution and the complex economies of content peering. Bell Labs Technical Journal, 17 (1):163–184, June 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lalli:2014:NSJ

[Lal14]  Roberto Lalli.  A new

scientific journal takes the scene: The birth of Reviews of Modern Physics. Annalen der Physik 8 (Berlin, Germany), 526(9–10):A83– A87, October 2014. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL http://

                                        onlinelibrary.wiley.com/
                                        enhanced/doi/10.1002/andp.
                                        201400810/.

Leost:2012:IRC

[LARJ12] Yann Léost, Moussa Abdi, Robert Richter, and Michael Jeschke. Interference rejection combining in LTE networks. Bell Labs Technical Journal, 17(1):25–49, June 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Li:2012:DPC

[LBC12] Xiang Yang Li, Kim Brouard, and Yigang Cai. Dynamic policy and charging control framework. Bell Labs Technical Journal, 17(1): 105–123, June 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Liu:2010:MFN

[LBTC10] Xiang Liu, Fred Buchali, Robert W. Tkach, and Sethumadhavan Chandrasekhar. Mitigation of fiber nonlinear impairments in polarization division multiplexed OFDM transmission. Bell Labs Technical Journal, 14 (4):47–59, Winter 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lou:2012:PVW

[LBW^{+}12] Zhe Lou, Jan Bouwen, Koen Willaert, Sigurd Van Broeck, Marc Van den Broeck, Senka Zubic, Wolfgang Van Raemdonck, Hans Van Herreweghe, and Dennis Dams. PresenceScape: Virtual world mediated rich communication. Bell Labs Technical Journal, 16(4):219–242, March 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ling:2013:CGH

[LCCV13] Jonathan Ling, Dmitry Chizhik, Chung Shue Chen, and Reinaldo A. Valenzuela. Capacity growth of heterogeneous cellular networks. Bell Labs Technical Journal, 18 (1):27–40, June 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lee:2010:ACI

[Lee10]  Anne Y. Lee.  Applica-

tion creation for IMS systems through macro-enablers and Web 2.0 technologies. Bell Labs Technical Journal, 15 (1):23–51, June 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lehoux:2013:IVC

[Leh13] Natalie Lehoux. Interaction that values co-creation in the design of services. Bell Labs Technical Journal, 17(4):145– 156, March 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lei:2010:ECL

[LH10] Bo Lei and Harry Hofmann. Energy consumption and low power design of optical equipment. Bell Labs Technical Journal, 15(2):169– 174, September 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Liu:2010:FRD

[Liu10] Lydia He Liu. The Freudian robot: digital media and the future of the unconscious. University of Chicago Press, Chicago, IL, USA, 2010. ISBN 0-226-48682-6 (hardcover), 0-226-48683-4 (paperback). xi + 302 pp. LCCN Q325 .L58 2010.

Lavigne:2013:SDT

[LLB^{+}13] Bruno Lavigne, Mathieu Lefrançois, Elodie Balmefrezol, Christine Bresson, Francesco Vacondio, Jean-Christophe Antona, Emmanuel Seve, and Olivier Rival. System design tool for high bit rate terrestrial transmission systems with coherent detection. Bell Labs Technical Journal, 18(3):251– 266, December 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Liem:2011:MCC

[LM11] Michael Liem and Veena B.

Mendiratta. Mission critical communication networks for railways. Bell Labs Technical Journal, 16(3):29–46, December 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Luton:2011:SMT

[LMMT11] Grégoire Luton, Rani Makké, Jens Mueckenheim, and Pascal Treillard. Support of mobile TV over an HSPA network. Bell Labs Technical Journal, 16(2):153– 168, September 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Layec:2013:EON

[LMV^{+}13] Patricia Layec, Annalisa Morea, Francesco Vacondio, Olivier Rival, and Jean-Christophe Antona. Elastic optical networks: The global evolution to software configurable optical networks. Bell Labs Technical Journal, 18 (3):133–151, December 2013. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Li:2013:HPP

[LT13] Frank Li and Alan Tang. High performance payload conversion method. Bell Labs Technical Journal, 18(2):135– 142, September 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Leven:2013:ONT

[LV13] Andreas Leven and Eve L. Varma. Optical networking and technology innovation. Bell Labs Technical Journal, 18(3):1–3, December 2013. CODEN BLTJFD.

ISSN 1089-7089 (print), 1538-7305 (electronic).

Lee:2011:O

[LW11] Anne Lee and Thomas L.

Wood. Overview. Bell Labs Technical Journal, 15 (4):1–3, March 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Mampaey:2011:SRT

[Mam11] Marcel J. Mampaey. Secure remittance transaction to bankless consumers in a fragmented applications market. Bell Labs Technical Journal, 16(2):219–233, September 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Marzetta:2015:MMI

[Mar15] Thomas L. Marzetta. Massive MIMO: An introduction. Bell Labs Technical Journal, 20(1):11–22, ???? 2015. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

McGarry:2011:BIT

[McG11] Patrick M. McGarry. Built it and they didn’t come: The right way to build and maintain developer communities. Bell Labs Technical Journal, 16(2):251–262, September 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Mirowski:2014:DFS

[MCHY14] Piotr Mirowski, Sining Chen, Tin Kam Ho, and Chun-Nam Yu. Demand forecasting in Smart grids. Bell Labs Technical Journal, 18(4):135– 158, March 2014. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

McGee:2012:PSN

[MCP12] Andrew R. McGee, Matthieu Coutière, and Maria E. Palamara. Public safety network security considerations. Bell Labs Technical Journal, 17(3):79–86, December 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Menten:2011:NEM

[MCS11] Lawrence E. Menten, Aiyou Chen, and Dimitrios Stiliadis. NoBot: Embedded malware detection for endpoint devices. Bell Labs Technical Journal, 16(1):155–170, June 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Molina-Garcia:2013:EBF

[MGCSA^{+}13] Mariano Molina-García, Jaime Calle-Sánchez, José I. Alonso, Alfonso Fernández-Durán, and Félix Barba Barba. Enhanced in-building fingerprint positioning using femtocell networks. Bell Labs Technical Journal, 18(2):195– 211, September 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Marilly:2013:GIV

[MGMP13] Emmanuel Marilly, Arnaud Gonguet, Olivier Martinot, and Frédérique Pain. Gesture interactions with video: From algorithms to user evaluation. Bell Labs Technical Journal, 17(4):103–118, March 2013. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Maes:2010:ADO

[MGP10] Jochen Maes, Mamoun Guenach, and Michael Peeters. Autonomous dynamic optimization for digital subscriber line networks. Bell Labs Technical Journal, 15 (3):119–129, December 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Miller:2012:MVD

[MHP^{+}12] Raymond B. Miller, John P.

Hearn, Cezary Purzynski, Fernando Cuervo, and Michael Scheutzow. Mobile video delivery using network aware transcoding in an LTE network. Bell Labs Technical Journal, 16(4):43–61, March 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Milioris:2014:JSC

[MJ14] Dimitrios Milioris and Philippe Jacquet. Joint sequence complexity analysis: Application to social networks information flow. Bell Labs Technical Journal, 18(4):75–88, March 2014. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Morea:2010:WLR

[MKVD10] Annalisa Morea, Daniel C. Kilper, Dominique Verchere, and Richard Douville. Wavelength layer recovery in transparent optical networks. Bell Labs Technical Journal, 14 (4):193–211, Winter 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

McBride:2012:ASG

[MM12] Alan J. McBride and Andrew R. McGee. Assessing smart grid security. Bell Labs Technical Journal, 17(3):87– 103, December 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Mheich:2012:OUS

[MMD12] Zeina Mheich, Marie-Line Alberi Morel, and Pierre Duhamel. Optimization of unicast services transmission for broadcast channels in practical situations. Bell Labs Technical Journal, 17 (1):5–23, June 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Maison:2013:SBJ

[MMDZ13] Rafal Maison, Ewelina Majda, Andrzej P. Dobrowolski, and Maciej Zakrzewicz. Similarity based join over Audio feeds in a multimedia data stream management system. Bell Labs Technical Journal, 18(1):195–212, June 2013. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Matthews:2010:PEE

[MMN^{+}10] H. Scott Matthews, Thomas B.

Morawski, Amy L. Nagengast, Gerard P. O’Reilly, David D. Picklesimer, Raymond A. Sackett, and Paul P. Wu. Planning energy-efficient and eco-sustainable telecommunications networks. Bell Labs Technical Journal, 15 (1):215–236, June 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Mirowski:2014:PRF

[MMWH14] Piotr Mirowski, Dimitrios Milioris, Philip Whiting, and Tin Kam Ho. Probabilistic radio-frequency fingerprinting and localization on the run. Bell Labs Technical Journal, 18(4):111– 133, March 2014. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Maes:2015:PPF

[MN15] Jochen Maes and Carl J. Nuzman. The past, present, and future of copper access. Bell Labs Technical Journal, 20 (1):1–10, ???? 2015. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Matthews:2010:OGM

[MNPW10] Elissa P. Matthews, Janet Nici, Beth Polonsky, and Joseph F. Wisniewski. Operating in the green: Modeling eco-friendly telecom network management. Bell Labs Technical Journal, 15(2):175– 192, September 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Macias:2013:OSH

[MPR13] Jesús Macias, Pablo Pérez, and Jaime J. Ruiz. Optimal size of an HDTV set during the transition period from SDTV to HDTV. Bell Labs Technical Journal, 17(4):119– 131, March 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Morea:2013:PES

[MPS^{+}13] Annalisa Morea, Jordi Perelló, Salvatore Spadaro, Dominique Verchère, and Martin Vigoureux. Protocol enhancements for ‘greening’ optical networks. Bell Labs Technical Journal, 18(3):211–230, December 2013. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

McKiou:2011:LBS

[MS11] Kevin W. McKiou and Arun Sankisa. Location based service extensions for general communications and application enablement. Bell Labs Technical Journal, 16(2):39– 56, September 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Marilly:2011:CBA

[MSA^{+}11] Emmanuel Marilly, Christophe Senot, Xavier Andrieu, Bertrand Boidart, Armen Aghasaryan, and Alexis Germaneau. Communitybased applications. Bell Labs Technical Journal, 15 (4):93–109, March 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Maison:2011:MHP

[MSZB11] Rafal Maison, Daryl J.

Steen, Maciej Zakrzewicz, and Zenon Biniek. Monitoring high performance data streams in vertical markets: Theory and applications in public safety and healthcare. Bell Labs Technical Journal, 16(3):163–179, December 2011. CODEN BLTJFD.

ISSN 1089-7089 (print), 1538-7305 (electronic).

Mullender:2012:PCC

[Mul12] Sape J. Mullender. Predictable Cloud computing. Bell Labs Technical Journal, 17(2):25–39, September 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Mampaey:2012:NCD

[MV12] Marcel Mampaey and Álvaro Nunez Villegas. A network-centric DRM for online scenarios. Bell Labs Technical Journal, 17(3):129–133, December 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Mullender:2012:CCN

[MW12] Sape J. Mullender and Thomas L. Wood. Cloud computing: New opportunities for telecom providers. Bell Labs Technical Journal, 17(2):1–3, September 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Maison:2011:PBL

[MZ11] Rafal Maison and Maciej Zakrzewicz. Prediction-based load shedding for burst data streams. Bell Labs Technical Journal, 16(1):121–132, June 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Maggio:2013:FSO

[MZZ13] Santo Maggio, Qingnan Zhang, and Zhongxu Zhao. Free-space optical backplane prototype for telecommunication equipment in the petabit/s range. Bell Labs Technical Journal, 18(3):285– 291, December 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Nithi:2011:SPM

[NdLvW11] Nachi K. Nithi and Adriaan J. de Lind van Wijngaarden. Smart power management for mobile handsets. Bell Labs Technical Journal, 15(4):149–168, March 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Novo:2013:MIR

[NDZ13] Robert A. Novo, Christopher J. Davolos, and Z. John Zhao. Measuring the impact of redirecting and offloading mobile data traffic. Bell Labs Technical Journal, 18(1):81–103, June 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Noll:2011:BLM

[NG11] A. Michael Noll and Michael N.

Geselowitz, editors. Bell Labs memoirs: voices of innovation. IEEE History Center, Lexington, KY, USA, 2011.

ISBN 1-4636-7797-9. v + 372 pp. LCCN TK415.B45 B35 2011.

Newbury:2011:CEI

[NGYG11] Mark E. Newbury, Robert E. Goldstein, Hong Yang, and Asif D. Gandhi. Cost effective improvement of EV-DO user experience through contention-based carrier addition. Bell Labs Technical Journal, 16(1):171–185, June 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Orlandi:2011:ICE

[OBT^{+}11] Barbara Orlandi, Erick Bizouarn, François Taburet, Frank Scahill, Jean-Luc Lafragette, Richard Evenden, Simon Ringland, Stephen Johnson, Telemaco Melia, Tim Twell, and Yacine El Mghazli. Improving the customer experience for heterogeneous wireless access. Bell Labs Technical Journal, 15 (4):23–44, March 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Payette:2011:ERE

[PCSS11] Chuck Payette, Bruce Cilli, Zulfiquar Sayeed, and Sameer Sharma. Enabling reliable and efficient wireless video uplink. Bell Labs Technical Journal, 16(2):169– 191, September 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Pohlmann:2013:LCT

[PDP^{+}13] Wolfgang Pöhlmann, Bernhard Deppisch, Thomas Pfeiffer, Carlo Ferrari, Mark Earnshaw, Alex Duque, Robert L. Farah, Joe Galaro, James Kotch, Man Fai Lau, Doutje T. van Veen, and Peter Vetter. Low cost TWDM by wavelength-set division multiplexing. Bell Labs Technical Journal, 18(3):173– 193, December 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Paparella:2013:AAE

[PG13] Andrea Paparella and Riccardo Gemelli. An analytical approach for estimating optimal multi-layer network cost, exploiting the flexiblegrid optical technology. Bell Labs Technical Journal, 18 (3):231–250, December 2013. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Pain:2013:HSU

[PL13] Frédérique Pain and Scott Lively. Human science and user experience. Bell Labs Technical Journal, 17(4): 1–4, March 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Perez:2011:EPL

[PMRG11] Pablo Pérez, Jesús Macías, Jaime J. Ruiz, and Narciso García. Effect of packet loss in video quality of experience. Bell Labs Technical Journal, 16(1):91–104, June 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Perez:2012:MVV

[PRV^{+}12] Pablo Pérez, Jaime J. Ruiz, Alvaro Villegas, Karel Van Damme, Christian Van Boven, Joel Dupont, and Pedro A. Molina Salmerón. Multi-vendor video headend convergence solution. Bell Labs Technical Journal, 17 (1):185–200, June 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Phadke:2013:PSC

[PUM^{+}13] Chitra Phadke, Huseyin Uzunalioglu, Veena B. Mendiratta, Dan Kushnir, and Derek Doran. Prediction of subscriber churn using social network analysis. Bell Labs Technical Journal, 17 (4):63–75, March 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Panshenskov:2011:GMN

[PV11] Mikhail A. Panshenskov and Alexander T. Vakhitov. General method for network providers to choose applications with guaranteed QoS. Bell Labs Technical Journal, 16(2):57–61, September 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Rader:2011:OSS

[Rad11] Reinhard Rader. Operations support systems for mission critical public safety communication networks. Bell Labs Technical Journal, 16(3):151– 162, December 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Radhakrishnan:2012:AAS

[Rad12] Ganesan Radhakrishnan. Adaptive application scaling for improving fault-tolerance and availability in the Cloud. Bell Labs Technical Journal, 17 (2):5–14, September 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Renaudier:2010:TLH

[RBPC^{+}10] Jérémie Renaudier, Oriol Bertran-Pardo, Gabriel Charlet, Massimiliano Salsi, Haïk Mardoyan, Patrice Tran, and Sébastien Bigo. 8 Tb/s long haul transmission over low dispersion fibers using 100 Gb/s PDM-QPSK channels paired with coherent detection. Bell Labs Technical Journal, 14(4):27–45, Winter 2010. CODEN BLTJFD.

ISSN 1089-7089 (print), 1538-7305 (electronic).

Robinson:2012:SVQ

[RJC12] David C. Robinson, Yves Jutras, and Viorel Craciun. Subjective video quality assessment of HTTP adaptive streaming technologies. Bell Labs Technical Journal, 16 (4):5–23, March 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Razavi:2010:FRL

[RKC10] Rouzbeh Razavi, Siegfried Klein, and Holger Claussen. A fuzzy reinforcement learning approach for self-optimization of coverage in LTE networks. Bell Labs Technical Journal, 15(3):153–175, December 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Rottondi:2013:ORM

[RTG13] Cristina Rottondi, Massimo Tornatore, and Giancarlo Gavioli. Optical ring metro networks with flexible grid and distance-adaptive optical coherent transceivers. Bell Labs Technical Journal, 18 (3):95–110, December 2013. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Raybon:2010:GDF

[RWS^{+}10] Gregory Raybon, Peter J. Winzer, Haoyu Song, Andrew Adamiecki, Stephen Corteselli, Alan H. Gnauck, Thomas Kissell, Daniel A. Fishman, Nat M. Denkin, Yuan-Hua Kao, Terry L. Downs, Anthony Carenza, Stephan Scrudato, Edward H. Goode, William A. Thompson, Christopher R. Doerr, Lawrence L. Buhl, Tiejun Xia, Glenn Wellbrock, Wang Lee, Greg Lyons, Peter Hofmann, Tina T. Fisk, E. Bert Basch, William J. Kluge, Johnny R. Gatewood, Tetsuya Kawanishi, Kaoru Higuma, and Yves Painchaud. 100 Gb/s DQPSK field trial: Live video transmission over an operating LambdaXtreme(R) network. Bell Labs Technical Journal, 14(4):85–113, Winter 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Saniee:2015:SAL

[San15] Iraj Saniee. Scalable algorithms for large and dynamic networks: Reducing big data for small computations. Bell Labs Technical Journal, 20 (1):23–33, ???? 2015. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Saniee:2010:O

[SB10] Iraj Saniee and Ulrich Barth.

Overview. Bell Labs Technical Journal, 15(3):1–4, December 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sabnis:2012:OCS

[SC12] Suhasini Sabnis and Doug Charles. Opportunities and challenges: Security in eHealth. Bell Labs Technical Journal, 17(3):105– 111, December 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Salsi:2010:RAS

[SCR^{+}10] Massimiliano Salsi, Gabriel Charlet, Jérémie Renaudier, Oriol Bertran-Pardo, Haïk Mardoyan, Sébastien Bigo, Philippe Plantady, and Alain Calsat. Recent advances in submarine optical communication systems. Bell Labs Technical Journal, 14(4):131– 147, Winter 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Schmalen:2013:FEC

[SdLvWtB13] Laurent Schmalen, Adriaan J. de Lind van Wijngaarden, and Stephan ten Brink. Forward error correction in optical core and optical access networks. Bell Labs Technical Journal, 18(3):39–66, December 2013. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sharp:2012:LCB

[Sha12]  Ron Sharp.  Latency in

Cloud-based interactive streaming content. Bell Labs Technical Journal, 17(2):67– 80, September 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Scanlon:2010:FEA

[SKL10] Patricia Scanlon, Irwin O.

Kennedy, and Yongheng Liu. Feature extraction approaches to RF fingerprinting for device identification in femtocells. Bell Labs Technical Journal, 15(3):141– 151, December 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Saniee:2017:WPG

[SKPW17] Iraj Saniee, Sanjay Kamat, Subra Prakash, and Marcus Weldon. Will productivity growth return in the new digital era. Bell Labs Technical Journal, 22(??):1–18, ???? 2017. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic). URL

https://ieeexplore.ieee.
org/document/7951155/.

Schoinianakis:2018:QVQ

[SML18] Dimitrios Schoinianakis and Enrique Martín-López. Quo vadis qubit? Bell Labs Technical Journal, 23(??): 1–17, ???? 2018. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic). URL

https://ieeexplore.ieee.
org/document/8526353/.

Schwan:2011:FCB

[SSG^{+}11] Nico Schwan, Klaus Satzke, David Griffin, Bertrand Mathieu, and Toufik Ahmed. Fostering cooperation between overlay applications and underlying networks. Bell Labs Technical Journal, 16(2):193–198, September 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sahm:2013:OAF

[SSM^{+}13] Hans Sahm, Matthias Sauppe, Erik Markert, Thomas Horn, Ulrich Heinkel, and Klaus-Holger Otto. Optimized ASIC/FPGA design flow for energy efficient network nodes. Bell Labs Technical Journal, 18(3):195– 209, December 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Spiess:2014:UBD

[STD^{+}14] Jeffrey Spiess, Yves T’Joens, Raluca Dragnea, Peter Spencer, and Laurent Philippart. Using big data to improve customer experience and business performance. Bell Labs Technical Journal, 18(4): 3–17, March 2014. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sabnis:2012:ISN

[SVHM12] Suhasini Sabnis, Marc Verbruggen, John Hickey, and Alan J. McBride. Intrinsically secure next-generation networks. Bell Labs Technical Journal, 17(3):17–36, December 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Stanze:2013:HNL

[SW13] Oliver Stanze and Andreas Weber. Heterogeneous networks with LTE-advanced technologies. Bell Labs Technical Journal, 18(1): 41–58, June 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Theeten:2014:TOP

[TBC^{+}14] Bart Theeten, Ivan Bedini, Peter Cogan, Alessandra Sala, and Tommaso Cucinotta. Towards the optimization of a parallel streaming engine for telco applications. Bell Labs Technical Journal, 18(4):181– 197, March 2014. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Thottan:2019:NCG

[TDK^{+}19] M. Thottan, C. Di Martino, Y. Kim, G. Atkinson, N. Choi, N. Mohanasamy, L. Jagadeesan, V. Mendiratta, J. E. Simsarian, and B. Kozicki. The network OS: Carrier-grade SDN control of multi-domain, multi-layer networks. Bell Labs Technical Journal, 24:1–26, December 2019. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Tamm:2010:ESS

[THR10] Oliver Tamm, Christian Hermsmeyer, and Allen M. Rush. Eco-sustainable system and network architectures for future transport networks. Bell Labs Technical Journal, 14(4):311–327, Winter 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Thiebaut:2012:UTW

[THU12] Laurent Thiébaut, Wim Henderickx, and Alistair Urie. Using a trusted WLAN network to offload mobile traffic and leverage deployed broadband network gateways. Bell Labs Technical Journal, 17 (1):89–104, June 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Tkach:2010:SOC

[Tka10] Robert W. Tkach. Scaling optical communications for the next decade and beyond. Bell Labs Technical Journal, 14 (4):3–9, Winter 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Tang:2010:FIR

[TQ10] Xiutao Tang and Cuilu Qu. Facial image recognition based on fractal image encoding. Bell Labs Technical Journal, 15(1):209–214, June 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Timmers:2013:TCA

[TVV^{+}13] Michael Timmers, Dirk Vanderhaegen, Danny Van Bruyssel, Mamoun Guenach, and Jochen Maes. Transmitter-controlled adaptive modulation: Concepts and results. Bell Labs Technical Journal, 18(1): 153–169, June 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Urie:2013:EMB

[URRH13] Alistair Urie, Ashok N.

Rudrapatna, Chandrasekharan Raman, and Jean-Marc Hanriot. Evolved multimedia broadcast multicast service in LTE: An assessment of system performance under realistic radio network engineering conditions. Bell Labs Technical Journal, 18(2):57– 76, September 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Urie:2013:O

[UV13] Alistair Urie and Reinaldo Valenzuela. Overview. Bell Labs Technical Journal, 18 (2):1–3, September 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Verdot:2011:APW

[VBB11] Vincent Verdot, Gérard Burnside, and Nicolas Bouché. An adaptable and personalized Web telecommunication model. Bell Labs Technical Journal, 16(1):3–17, June 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Viswanathan:2013:MVO

[VDB^{+}13] Harish Viswanathan, Danny De Vleeschauwer, Andre Beck, Steven Benno, Raymond B. Miller, Gang Li, Mark M. Clougherty, and David C. Robinson. Mobile video optimization at the base station: Adaptive guaranteed bit rate for HTTP adaptive streaming. Bell Labs Technical Journal, 18(2):159– 174, September 2013. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Verhoeyen:2012:OVS

[VDD12] Marc Verhoeyen, Johan De Vriendt, and Danny De Vleeschauwer. Optimizing for video storage networking with recommender systems. Bell Labs Technical Journal, 16(4):97–113, March 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Vidal:2013:RAI

[VGF13] Guillermo Martin Vidal, Marjan Geerts, and Mohamed Ali Feki. The role of affordances and interaction bits in the design of a new tangible programming interface: a preliminary result. Bell Labs Technical Journal, 17(4):157–174, March 2013. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

VanBroeck:2011:RTV

[VLV11] Sigurd Van Broeck, Zhe Lou, and Marc Van den Broeck. Real-time 3D video communication in 3D virtual worlds: Technical realization of a new communication concept. Bell Labs Technical Journal, 16 (1):63–76, June 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Villegas:2012:NAC

[VPC^{+}12] Álvaro Villegas, Pablo Pérez, José María Cubero, Enrique Estalayo, and Narciso García. Network assisted content protection architectures for a connected world. Bell Labs Technical Journal, 16 (4):85–96, March 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

VandeVoorde:2012:I

[VV12] Ingrid Van de Voorde and Willem Verbiest. Introduction. Bell Labs Technical Journal, 16(4):1–3, March 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Vieira:2012:CVP

[VVG12] Elisangela Rodrigues Vieira, Dario Vieira, and Jean-Pierre Gallois. Challenges of a validation process based on models: An industrial case study. Bell Labs Technical Journal, 17(1):229–246, June 2012. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Viswanathan:2014:PPF

[VW14] H. Viswanathan and M. Weldon. The past, present, and future of mobile communications. Bell Labs Technical Journal, 19(1):8–21, ???? 2014. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Walid:2010:SAD

[WC10] Anwar Walid and Aiyou Chen. Self-adaptive dynamic bandwidth allocation for GPON. Bell Labs Technical Journal, 15(3):131– 139, December 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Winzer:2014:SMF

[Win14] P. J. Winzer. Spatial multiplexing in fiber optics: The 10× scaling of metro/core capacities. Bell Labs Technical Journal, 19(1):22–30, ???? 2014. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Wood:2011:I

[WM11] Thomas H. Wood and Olivier Martinot. Introduction. Bell Labs Technical Journal, 16 (1):1–2, June 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Xie:2011:ULS

[XW11] Junqing Xie and Shuai Wang.

A unified location sharing service with end user privacy control. Bell Labs Technical Journal, 16(2):5–20, September 2011. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Xie:2010:OPM

[XWH^{+}10] Chongjin Xie, Dieter Werner, Herbert Haunstein, Robert M. Jopson, Sethumadhavan Chandrasekhar, Xiang Liu, Yan Shi, Siegfried Gronbach, Thomas Link, and Konrad Czotscher. Optical polarization mode dispersion compensators for high-speed optical communication systems.

Bell Labs Technical Journal, 14(4):115–129, Winter 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Yang:2012:DFM

[Yan12]  Hong Yang.  Differential

fade margin and handoff gain in wireless cellular networks. Bell Labs Technical Journal, 17(1):51–65, June 2012. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Yassine:2011:FCS

[Yas11] Abdulsalam Yassine. Future communication services: Application enablement and the role of software agent architectures. Bell Labs Technical Journal, 16(2):271– 276, September 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Yu:2011:SGC

[YJ11] Qin Yu and Raymond J. Johnson. Smart grid communications equipment: EMI, safety, and environmental compliance testing considerations. Bell Labs Technical Journal, 16(3):109– 131, December 2011. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Zami:2013:CFF

[Zam13] Thierry Zami. Current and future flexible wavelength routing cross-connects. Bell Labs Technical Journal, 18 (3):23–38, December 2013. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Zami:2010:NMP

[ZML10] Thierry Zami, Annalisa Morea, and Florence Leplingard. A new method to plan more realistic optical transparent networks. Bell Labs Technical Journal, 14(4):213– 226, Winter 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Zhao:2010:SVD

[ZYG10] Shengjie Zhao, Mingli You, and Luoning Gui. Scalable video delivery over MIMO OFDM wireless systems using joint power allocation and antenna selection. Bell Labs Technical Journal, 15 (1):133–148, June 2010. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).